### v1

# Omni-Think: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

Abstract
The advancement of general-purpose artificial intelligence relies on large language models (LLMs) that excel across a wide range of tasks, from structured reasoning to creative generation. However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning. In this work, we introduce Omni-Think, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations. Our approach enables consistent optimization across task types and scales RL-based training to subjective domains. We further investigate training strategies, demonstrating that a curriculum-based progression that orders tasks from structured to open-ended improves performance and reduces forgetting. Experimental results across four domains reveal that curriculum learning improves performance by 5.2% over joint training and 9.1% over model merging. These results highlight the importance of task-aware sampling and hybrid supervision in scaling RL-based post-training for general-purpose LLMs.
1 Introduction
Figure 1: OMNI-THINK Framework Overview. A unified multi-task RL framework for training a single LLM across structured and open-ended tasks. (1) Backward Transfer quantifies task forgettability and informs (2) a curriculum that schedules tasks from least to most forgettable. (3) Sampled tasks construct prompts. (4) The LLM generates completions, which are (5) scored by hybrid rewards: verifiable (e.g., test case execution), generative (e.g., LLM-as-a-Judge), and auxiliary (e.g., format adherence), among others. (6) The policy is updated via Multi-Task GRPO. The process repeats via iterative updates (bottom loop).
As Large Language Models (LLMs) (Hurst et al., 2024; Liu et al., 2024; Dubey et al., 2024; Yang et al., 2024) evolve into general-purpose agents with applications ranging from creative writing (Marco et al., 2025) to robotics (Gemini-Robotics-Team et al., 2025), the challenge of effective post-training becomes increasingly critical. While Supervised Fine-Tuning (SFT) is the dominant paradigm for adapting LLMs to downstream tasks, it often encourages memorization rather than robust generalization, especially on tasks that differ from the training distribution or require flexible, context-dependent reasoning (Chu et al., 2025).
Reinforcement Learning (RL) has emerged as a promising path to improve generalization, especially in structured tasks like math and coding, where verifiable, rule-based rewards can be effective (DeepSeek-AI et al., 2025; Luo et al., 2025; Kimi-Team et al., 2025). Methods such as Group Relative Policy Optimization (GRPO) have shown that even coarse signals can guide LLMs toward structured, chain-of-thought responses (Shao et al., 2024). However, most RL methods are tailored to deterministic, easily verifiable tasks, limiting their utility in open-ended settings such as question answering and creative writing. Moreover, training LLMs across multiple tasks remains challenging because it requires optimizing for diverse forms of feedback signals, including binary correctness checks in structured tasks and subjective, preference-based judgments in generative ones.
To address this, we propose Omni-Think, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks. Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks. We further show that curriculum training, which sequences tasks from structured to open-ended based on Backward Transfer (BWT) measurements of task forgettability, significantly improves generalization and reduces forgetting. Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
Our key contributions are: (1) We propose Omni-Think, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains. (2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains. (3) We demonstrate that not all multi-task training strategies are equally effective. Curriculum-based task scheduling outperforms uniform sampling and improves generalization in multi-task RL.
2 Related Work
Large Language Models and Multi-Task Learning
Early work like (Sanh et al., 2021) showed that multi-task prompted training can encourage zero-shot generalization. Dong et al. (2023) further analyzed how mixing SFT data across domains can cause performance conflicts and forgetting, proposing Dual-stage Mixed Fine-tuning to alleviate these effects. However purely supervised objectives often encourage memorization rather than transferable reasoning. The Qwen3 model series (Yang et al., 2025) employs a four-stage post-training pipeline in the order of reasoning, non-reasoning, and general-domain under a mix of supervised fine-tuning and reinforcement learning. In comparison, the post-training process for Command-A (Cohere et al., 2025) alternates between training multiple expert models separately and merging the experts’ parameters into a “Soup Model” during its SFT and RL steps, before the model undergoes a polishing phase of preference alignment. In contrast, our work integrates multi-task learning directly into a single RL framework. Its backward-transfer-guided curriculum orders tasks from least to most forgettable, drawing on continual learning insights (Lopez-Paz & Ranzato, 2017) to reduce interference and maintain stable cross-task performance.
Large Language Models and Reinforcement Learning
Reinforcement Learning with Verified Rewards has demonstrated effectiveness for tasks with deterministic correctness signals such as math or code generation (Lambert et al., 2024; Shao et al., 2024; Kimi-Team et al., 2025; Guo et al., 2025). Recent frameworks like General-Reasoner (Ma et al., 2025), Nemotron-Crossthink (Akter et al., 2025) and X-REASONER (Liu et al., 2025) expand this to broader reasoning by blending multi-domain corpora and structured answer templates. However, these tasks still largely remain largely confined to verifiable STEM problems or multiple-choice formats, leaving open-ended generation, such as creative writing, insufficiently addressed. To bridge this gap, Su et al. (2025) propose a generative reward model (GRM) to replace rule-based signals. Although this improves RL and makes it applicable to general-domain QA when references exist, the approach is still restricted to verifiable tasks. In contrast, our approach integrates hybrid verifiable and preference-based rewards within a single RL loop, enabling consistent optimization across both structured and open-ended tasks. Moreover, our curriculum design, guided by backward transfer, helps maintain stable cross-task performance even for tasks lacking deterministic evaluation criteria.
3 Problem Formulation
We consider the problem of training a single language model policy to perform well across a diverse set of tasks, including mathematical reasoning, code generation, question answering, and creative writing. Let $\mathcal{T}=\{T_{1},\ldots,T_{K}\}$ be a collection of $K$ tasks, where each task $T_{k}$ has a dataset $\mathcal{D}_{k}$ of input-output pairs $(q_{k},o^{*}_{k})$, where $q_{k}\in\mathcal{Q}_{k}$ is an input prompt, and $o^{*}_{k}\in\mathcal{O}^{*}_{k}$ is the corresponding ground-truth output response.
The model, parameterized by $\theta$, defines a conditional distribution $\pi_{\theta}(y_{k}\mid x_{k})$. Supervised fine-tuning (SFT) aims to maximize the likelihood of ground-truth outputs by minimizing:
$\min_{\theta}\sum_{k=1}^{K}\mathbb{E}_{(x_{k},y_{k})\sim\mathcal{D}_{k}}\left[\sum_{t=1}^{T}-\log\pi_{\theta}(o^{*}_{k,t}\mid q_{k},o^{*}_{k,<t})\right],$ (1)
where $o^{*}_{k,t}$ denotes the $t$-th token of $o^{*}_{k}$, and $T$ is the number of tokens in the output.
While effective in-domain, SFT often overfits to training data and fails to generalize, particularly when task formats vary. Recent studies (Chu et al., 2025) show that reinforcement learning (RL) with outcome-based rewards better promotes cross-task generalization. Motivated by this, we adopt a multi-task RL (MTRL) formulation with task-specific reward functions $R_{k}$. The goal is to learn a unified policy $\pi_{\theta}$ that maximizes the expected reward over the task distribution:
$\max_{\theta}\;\mathcal{J}(\theta)=\mathbb{E}_{T_{k}\sim P(\mathcal{T})}\left[\mathbb{E}_{q_{k}\sim\mathcal{D}_{k},o_{k}\sim\pi_{\theta}(\cdot\mid x_{k})}\left[R_{k}(o_{k})\right]\right],$ (2)
where $P(\mathcal{T})$ is the task sampling distribution, which determines task exposure during training, and $R_{k}$ is the task-specific reward function. This objective trains a single policy, $\pi_{\theta}$, that balances performance across tasks, enabling knowledge transfer, mitigating negative transfer, and improving generalization across domains.
4 Methodology
4.1 Omni-Think Framework Overview
We introduce Omni-Think, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
Verifiable Supervision.
For tasks with objective correctness signals, such as symbolic math and code generation, we define binary rewards based on symbolic matches, test case results, or other deterministic evaluators depending on the tasks.
Short-Form Open-Ended Supervision.
For concise generative tasks with known answers (e.g., General question answering (QA)), we transform the questions into open-ended prompts and embed the distractor responses, which are plausible but incorrect candidate answers, within the context. Rather than labeling options, we instruct the model to reason with <think></think> and produce answers in <answer></answer> format. A response is assigned a binary reward based on whether it matches any reference answer via string comparison or set membership, promoting semantic grounding and discouraging pattern memorization.
Long-Form Open-Ended Supervision.
For subjective tasks lacking ground truth (e.g., dialogue, writing), we use an LLM-as-a-Judge  (Chen et al., 2025) to assign a scalar reward based on rubric-aligned pairwise preferences between candidate outputs. This enables learning in domains where symbolic correctness is insufficient or intractable.
Together, these strategies enable consistent RL training across structured and generative tasks within a unified optimization framework.
4.2 Multi-Task Group Relative Policy Optimization
We extend the GRPO (Guo et al., 2025) algorithm to the multi-task setting by jointly optimizing over task-specific reward signals and reference policies. For each input prompt $q_{k}$, GRPO samples a group of outputs $\{o_{k,1},o_{k,2},\cdots,o_{k,G}\}$ from the old policy $\pi_{\theta_{old}}$. At each generation step $t$, the partial output is denoted $o_{k,i,<t}$, where $o_{k,i,<t}=[o_{k,i,1},\dots,o_{k,i,t-1}]$, that is, the input concatenated with the sequence of generated tokens up to step $t-1$. A reward function $R_{k}(o,\phi_{k})$ scores each output,where $\phi_{k}$ denotes task-specific information required for the evaluation, such as reference answers, test cases, or comparison outputs required for evaluation. The policy $\pi_{\theta}$ is updated to maximize expected return while controlling divergence from a reference policy.
Let
$\mu_{k}=\text{mean}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},\quad\sigma_{k}=\text{std}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},$ (3)
we define the policy ratio $\rho_{k,i,t}$ and the normalized advantage estimate $\hat{A}_{k,i,t}$ as
$\rho_{k,i,t}=\frac{\pi_{\theta}(o_{k,i,t}\mid q_{k},o_{k,i,<t})}{\pi_{\theta_{\text{old}}}(o_{k,i,t}\mid q_{k},o_{k,i,<t})},$ $\quad\hat{A}_{k,i,t}=\frac{R_{k}(o_{k,i},\phi_{k})-\mu_{k}}{\sigma_{k}}.$ (4)
This allows us to write the MT-GRPO objective as
$\begin{split}&\mathcal{J}_{\text{MT-GRPO}}(\theta)=\mathbb{E}_{k\sim K,q_{k}\sim\mathcal{D}_{k},\{o_{k,i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{k})}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{k,i}|}\sum_{t=1}^{|o_{k,i}|}\left\{\min\left[\rho_{k,i,t}\hat{A}_{k,i},\text{clip}\left(\rho_{k,i,t},1{-}\epsilon,1{+}\epsilon\right)\hat{A}_{k,i}\right]-\beta_{k}\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (5)
where
$\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]=\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1.$ (6)
The clipping parameter $\epsilon$ stabilizes updates by keeping policy ratios within a bounded range, following the PPO approach (Schulman et al., 2017). The KL divergence term regularizes the new policy towards the reference policy $\pi_{\text{ref}}$, weighted by a task-specific coefficient $\beta_{k}$.
4.3 Task Scheduling
We consider two strategies for optimizing a shared policy across multiple tasks: curriculum learning and joint training. Both utilize task-specific rewards but differ in scheduling. To guide task ordering, we adapt a Backward Transfer (BWT) metric to quantify forgetting in our multi-task setting. Following Lopez-Paz & Ranzato (2017), we define $BWT_{j}=P_{after,j}-P_{base,j}$, where $P_{base,j}$ is the performance on task $j$ from the original base model and $P_{after,j}$ is the performance on task $j$ after training on subsequent tasks. A negative $BWT_{j}$ indicates that performance on task $j$ degrades due to learning new tasks, measuring the extent of forgetting.
Curriculum Learning:
Prior work in curriculum learning prioritizes task complexity (e.g., easy to hard) for more effective learning (Bengio et al., 2009; Parashar et al., 2025). We instead adopt a forgetting-based curriculum that orders tasks according to their average forgettability in multi-task learning. Tasks with the least negative or even positive backward transfer are introduced earlier, serving as stable foundations for subsequent tasks that are more prone to catastrophic forgetting. In our setup, this coincides with increasing supervision subjectivity. This forgetting-based ordering reduces overall forgetting and improves generalization by minimizing destructive interference across tasks.
Joint Training:
All tasks are sampled concurrently from a shared distribution. While this promotes exposure, it introduces three key challenges: (1) simpler tasks may dominate without adaptive sampling due to the tendency of reward hacking in GRPO; (2) reward delays vary, requiring asynchronous updates; and (3) Kullback–Leibler (KL) regularization demands tuning per task due to varying sensitivities.
In both regimes, we apply MT-GRPO with normalized advantages and task-specific KL coefficients. At each step, a task $T_{k}$ is selected, completions are generated by the current policy $\pi_{\theta}$, and updates are applied based on the corresponding reward signal.
5 Experimental Setup
5.1 Training Datasets
We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation. For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning. We further subsample 12,000 examples to fit our compute budget. For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed. Each entry includes a code prompt and JSON-formatted unit tests for automatic validation. For General QA, we extract 5,500 examples from SuperGPQA (M-A-P, 2025), downsampled proportionally by question category. Each sample comprises a factual question paired with a plain-text answer. The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on one-turn completions. Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
5.2 Baselines
We adopt Qwen2.5-7b-Instruct as the base model for all our experiments (Yang et al., 2024). Its robust instruction-following ability makes it a suitable candidate for subsequent reinforcement learning training on both reasoning tasks and more general open-domain QA, as it relies on the model’s capacity to comprehend and follow the given prompts effectively.
Supervised Fine-Tuning (SFT):
In order to have a meaningful comparison with GRPO, we adopt a similar self-sampled data curation and fine-tuning approach with Rejection sampling Fine-Tuning (Yuan et al., 2023). We first prompt the base model to generate 128 chain-of-thought responses for our training dataset to ensure we end up with at least one correct response for most queries, then filter them based on the same accuracy reward signals used in GRPO training. We then perform supervised fine-tuning on Qwen2.5-7b-Instruct using these self-distilled responses. This provides a strong on-policy learning baseline that incorporates explicit reasoning steps through self-distillation from the base model.
Model Merging:
We employ TIES-Merging  (Yadav et al., 2023b) as our model‑merging baseline. It’s a simple yet effective method designed specifically for the multi-task setting that takes into consideration the interference between parameters from models trained on individual tasks during the merging process. It has demonstrated superior performance in multi-task learning compared to linear and task arithmetic approaches (Yadav et al., 2023a). To begin with, we conduct single-task GRPO training using individual task datasets and collect the model weights of the best checkpoints with the help of a validation set for each training run. We then merge the four single-task models using a scaling value $\lambda=1$.
5.3 Evaluation
We assess performance in each domain using dedicated held-out benchmarks aligned with the task’s unique evaluation criteria. The detailed evaluation set is presented as follows
Math Reasoning:
we compute accuracy over seven datasets: AIME24 (MAA, 2024), AMC23 (MAA, 2023), Gaokao2023EN (Liao et al., 2024), MATH-500 (Hendrycks et al., 2021), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024).
Code Generation:
we measure coding ability via pass@1 on BigCodeBench (Complete-Full) (Zhuo et al., 2024) and LiveCodeBench (24Oct-25Jan) (Jain et al., 2024).
General QA:
we report exact-match accuracy using the MMLU-Pro benchmark (Wang et al., 2024).
Creative Writing:
we evaluate the creative writing task using the role-play and creative writing subcategory of MT‑Bench (Zheng et al., 2023), reporting win rate against a GPT-4 (pre-gen dated June 16, 2023) model.
6 Results and Discussion
Figure 2: Performance gains across four task domains, comparing our Multi-Task (MT) framework (Joint and Curriculum variants) against baselines including Supervised Fine-Tuning (SFT) and Model Merging. Curriculum MT achieves the strongest results, particularly in open-ended tasks, showing that controlling how multi-task learning is structured is crucial for effective generalization.
6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
We evaluate Omni-Think across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization. Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results. Table 2 further details how these gains vary by benchmarks.
In Math, Curriculum MT achieves the highest average performance at 59.6%, with the clearest gains on more complex reasoning tasks such as MinervaMath and OlympiadBench. These benchmarks benefit from strong rule-based reward signals and backward-transfer-guided task ordering. In contrast, datasets like AMC23 show minimal change because their relatively high baseline scores likely reflect smaller question sets and potential pretraining overlap rather than robust multi-step problem-solving.
In General QA, Curriculum MT again performs best (52.2%), followed by Model Merging (49.8%) and Joint Multi-Task GRPO (48.8%). These improvements are driven by our Short-Form Open-Ended Supervision strategy: instead of generating responses in a fully open-ended and unconstrained fashion, the model is trained to produce complete answer strings given a diverse set of candidate responses, enabling the effective application of verifiable reward through simple string matching when training general-domain tasks.
For Code Generation, Curriculum MT achieves 35.4%, slightly ahead of Model Merging. Notably, we only evaluate on the subset of LiveCodeBench(24Oct-25Jan) problems released after Qwen2.5’s data cutoff, which ensures that these are unseen test items. This setup highlights Curriculum MT’s significant generalization gains on novel problems, explaining the larger improvements on LiveCodeBench relative to static benchmarks like BigCodeBench, where data overlap is more likely.
In Creative Writing, the introduction of our Long-Form Open-Ended Supervision strategy, employing the LLM-as-a-Judge framework for pairwise evaluation, results in significant performance boosts (Curriculum-Guided at 84.2% and Joint MT at 83.00%), underscoring the advantage of our generative reward approach in subjective, open-ended tasks.
These results support our central hypothesis: The Omni-Think Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
Figure 3: Average backward transfer (BWT) per target task, showing how training single-task GRPO on all other tasks affects it on average. A negative BWT indicates forgetting; a positive value indicates positive transfer.
Figure 4: Models trained to generate full-text answers outperform those trained to select letter choices. This format promotes deeper semantic understanding over shallow pattern matching or guessing.
6.2 Training Order Matters: From Least to Most Forgettable
Our analysis highlights that task ordering is critical for robust multi-task reinforcement learning. Figure 4 summarizes the average backward transfer (BWT) each task receives when training single-task GRPO models on other tasks. Specifically, the x-axis of Figure 4 represents the target task. A negative average BWT value for a target task indicates that it tends to suffer forgetting or experience a decrease in performance when other tasks are trained before it; a positive value suggests that the target task tends to benefit or experience positive transfer. This high-level view reveals that Creative Writing is the most vulnerable task, as it experiences the most negative average BWT received, while Coding is the most resilient, showing a positive average BWT received.
For example, Coding benefits on average after training on another task (average BWT received $+1.06\%$, indicating a positive transfer). In contrast, Creative Writing is the most vulnerable to forgetting: it has the highest magnitude of negative average BWT received ($-4.57\%$), meaning it easily loses performance when other tasks are trained first.
Figure 5 shows the full BWT matrix, where x-axis consists of target tasks and y-axis represent source tasks that are being trained on. Each cell $(row_{i},col_{j})$ in the matrix represents the BWT from source task $j$ to target task $i$. A negative value indicates that training on source task $j$ causes forgetting in target task $i$, while a positive value suggests positive transfer. We observe that structured tasks like Coding and Math are both less forgetting conducive as sources and tend to receive more positive transfers as targets from training on other tasks.
Guided by this analysis, our curriculum trains tasks in a least-to-most forgettable order: Coding $\rightarrow$ Math $\rightarrow$ General QA $\rightarrow$ Creative Writing. This progression grounds the model in stable, verifiable tasks before introducing those more prone to interference and forgetting. As confirmed by Table 1, reversing this order significantly harms performance.
Overall, our findings suggest that BWT patterns can offer a practical, empirically informed guide for task sequencing, helping reduce negative transfer and improve generalization in multi-task RL for LLMs. Further work is needed to test this approach across a broader range of tasks and domains, including those that require logical reasoning over graph-structured data  (Zhou et al., 2024) and knowledge-base retrieval  (Dehghan et al., 2024), as well as more diverse open-ended domains such as mixed-initiative collaborative storytelling and co-creativity  (Kreminski et al., 2024).
6.3 Output Format Matters: Full-Text Answers Enhance Generalization
We examine how output format impacts generalization by comparing models trained to generate full-text answers versus selecting letter choices in multiple-choice QA (MCQ). Using GRPO, we train two single-task policies on the training set, one prompted to produce full-text final answers at the end of its chain-of-thought completions, and the other to output only letter choices (e.g., “A”, “B”, “C”). As shown in Figure 4, the model trained to output full-text answer achieves significantly better generalization when evaluated with free-form QA prompts on MMLU Pro (51% vs. 41%). While the letter-choice model slightly outperforms when evaluated strictly on MCQ prompts, the full-text model remains competitive across both prompt formats.
These results suggest that training with complete, semantically grounded answers encourages deeper reasoning, improving the model’s ability to generalize beyond the specific format seen during training. In contrast, letter-choice training risks overfitting to shallow pattern matching, reducing transferability to realistic QA settings that often require articulated responses.
7 Conclusion
We presented OMNI-THINK, a unified framework that extends LLMs to handle both structured and open-ended tasks within a single policy. By integrating reinforcement learning with both rule-based verifiable rewards and generative preference-based supervision, our method improves generalization while addressing challenges such as forgetting and task interference. Our findings show that effective multi-task LLM post-training depends not only on reward design but also on how tasks are sequenced and optimized together. By ordering tasks from structured to open-ended domains based on backward transfer, OMNI-THINK minimizes forgetting and achieves consistent gains across diverse reasoning and generative tasks. Overall, it takes a step toward more general-purpose LLMs that can learn effectively from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation within a single, unified post-training framework.
Appendix A Appendix
A.1 Reward Estimation
Omni-Think employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework. We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison. Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation. While some rewards (e.g., math and code) are strictly deterministic, others, such as LLM-as-a-Judge comparisons, are inherently stochastic but executed at low decoding temperature to ensure stable and consistent supervision. All reward functions are designed to be domain-aware and automatable, supporting scalable reinforcement learning across both structured and generative tasks.
Primary Rewards.
Each task employs a tailored correctness criterion:
  • Math:
    $r_{\texttt{math}}(o,a)=\mathbbm{1}\left\{\texttt{verify}_{\texttt{math}}(o_{\texttt{ans}},a)=\texttt{true}\right\}\,,$
    where $a$ denotes the ground-truth answer, and $o_{\texttt{ans}}$ is the predicted answer extracted from the <answer> tags in the model output $o$. Symbolic equivalence between $o_{\texttt{ans}}$ and $a$ is verified using a deterministic parser.
  • Code Generation:
[…]
    where the generated code $o_{\texttt{ans}}$ is executed in a sandboxed environment and evaluated against the unit tests defined by $x$; $\models$ indicates logical satisfaction.
  • General QA:
    $r_{\texttt{qa}}(o,a)=\mathbbm{1}\left\{o_{\texttt{ans}}=a\right\}$
    which returns 1 if the predicted answer matches the ground-truth string exactly.
  • Creative Writing:
    $r_{\text{writing}}(o,o_{ref})=\begin{cases}1.0&\text{if }o\succ o_{ref}\\ 0.5&\text{if }o\sim o_{ref}\\ 0.0&\text{if }o\prec o_{ref}\end{cases}$
    where $o\succ o_{ref}$ indicates that the model output $o$ is preferred over the reference $o_{ref}$, $o\prec o_{ref}$ means the reference is preferred, and $o\sim o_{ref}$ denotes a tie. Preferences are judged via pairwise comparison by a fixed LLM-as-a-Judge model.
Auxiliary Rewards.
To encourage structured outputs, we define formatting-based rewards shared across tasks:
$r_{\texttt{format}}(o)=\mathbbm{1}\left\{\texttt{tags\_valid}(o)\right\}$
$r_{\texttt{tags}}(o)=\frac{1}{4}\cdot\left|\texttt{tags\_present}(o)\right|$
Here, tags_valid ensures proper nesting of <think> and <answer> tags, while tags_present counts required structural markers.
Total Reward.
We define the total reward as a weighted sum over both primary and auxiliary reward components. Let $\mathcal{F}_{k}=\{r^{(1)}_{k},r^{(2)}_{k},\dots,r^{(m)}_{k}\}$ denote the set of reward functions associated with task $k$, where each $r^{(j)}_{k}$ measures a different aspect of correctness. Given a model output $o$ and its associated evaluation context $\phi_{k}$, the total reward is computed as:
$\mathcal{R}_{k}(o,\phi_{k})=\sum_{r\in\mathcal{F}_{k}}w_{r}\cdot r(o,\phi_{k}),$
where $w_{r}\in[0,1]$ denotes the task-specific weight for reward component $r$. If a component reward is undefined, e.g., due to malformed or unparsable output, it is omitted from the sum. Samples with no valid reward components are excluded from policy updates.
A.2 Results
Figure 5: Backward transfer (BWT) matrix. Each cell shows the BWT for a target task (columns) when training a single-task GRPO model on a source task (rows). Positive BWT indicates forgetting; negative indicates positive transfer.
Table 1: Ablation study comparing curriculum Learning (Curr) to its reversed task ordering (Reverse-Curr). Ordering tasks from structured to open-ended domains consistently improves performance across all, while reversing this order leads to degraded results, especially in open-ended domains.
Task Base Model Curr Reverse-Curr
Math 55.9 59.6 58.2
General QA 41.5 52.2 22.6
Code Generation 29.6 35.4 32.7
Creative Writing 74.2 84.2 75.0
Table 2: Performance across benchmarks. MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math). Model Merge applies TIES-merging across tasks. Bolded values mark the best per row. Domains include Math (7 sets), MMLU-Pro (9 categories), Coding (2 sets), and Creative Writing (MT-Bench).
Task Base Model ST: Coding ST: Math ST: QA ST: Writing SFT Model Merging Joint MT Curriculum MT
Math
aime24 15.33 15.33 14.67 14.00 13.33 16.67 14.67 19.33 14.67
amc23 62.50 60.00 60.00 62.50 59.00 62.00 65.50 64.00 65.50
gaokao2023en 72.99 74.29 76.10 74.03 75.58 74.29 74.81 76.62 77.14
math500 78.20 78.80 80.40 75.40 79.20 76.80 79.80 77.60 81.00
minerva_math 64.34 63.97 66.54 63.24 61.76 65.07 66.18 68.38 71.69
OlympiadBench 42.07 42.96 43.70 41.33 42.96 42.96 41.78 43.56 47.41
General QA
Biology 57.60 56.76 52.30 67.36 59.00 66.25 65.55 67.22 68.76
Business 33.46 39.04 25.60 58.68 32.95 48.16 59.82 49.81 47.53
Chemistry 35.78 31.80 27.30 47.70 38.34 44.08 42.49 42.05 50.71
Computer Science 53.66 48.05 50.24 55.12 51.95 53.66 53.90 58.78 59.27
Economics 42.65 49.17 38.74 62.91 44.91 59.60 61.97 56.75 62.09
Engineering 28.28 31.27 20.43 37.46 26.63 37.77 38.08 35.81 37.05
Health 46.70 46.21 45.23 50.98 47.19 45.72 52.69 50.73 57.09
History 37.27 33.33 34.65 47.24 38.58 33.86 47.24 43.31 45.67
Law 23.16 23.98 20.62 27.88 23.25 26.79 26.61 27.52 29.70
Math 55.37 52.63 50.41 59.29 56.25 57.36 58.25 59.22 61.21
Other 44.26 40.04 39.72 50.97 43.94 46.43 51.84 49.89 53.25
Philosophy 36.87 34.27 33.27 43.89 35.47 38.20 41.48 42.08 42.89
Physics 41.11 37.41 30.79 53.66 41.57 49.81 46.73 48.04 55.58
Psychology 50.88 51.50 45.36 60.15 51.75 59.02 59.40 59.27 61.78
Code Generation
BigCodeBench 46.49 50.35 46.66 47.10 46.84 44.47 48.07 47.19 49.47
LiveCodeBench 12.71 21.80 13.07 13.79 13.31 14.21 20.78 17.04 21.30
Creative Writing
MT-Bench (Writing) 74.16 71.60 74.16 63.00 78.33 44.17 67.50 83.00 84.17
A.3 Detailed Hyper-Parameters
We show the hyperparameters used in our training in Table3
Table 3: Training Hyperparameters for All Training Settings. MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math).
Hyperparameter Curr. MT Joint MT ST:Coding ST:Math ST:QA ST:Writing SFT
Model Configuration
   Max Prompt Length 1024 1024 1024 1024 1024 1024 -
   Max Response Length 3072 3072 3072 3072 3072 3072 -
Training Settings
   Train Batch Size 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 128
   Learning Rate 1e-6 1e-6 1e-6 1e-6 1e-6 1e-6 2.5e-6
   Learning Scheduler Constant Constant Constant Constant Constant Constant Cosine
   Optimizer AdamW AdamW AdamW AdamW AdamW AdamW AdamW
   Grad Clip 1.0 1.0 1.0 1.0 1.0 1.0 1.0
   Max Epoch 3 3 3 3 3 3 3
RL Settings
   KL Beta 0.00 0.02 0.001 0.04 0.04 0.00 -
   Clip Ratio Low 0.2 0.2 0.2 0.2 0.2 0.2 -
   Clip Ratio High 0.2 0.2 0.2 0.2 0.2 0.2 -
   $N$ Rollouts 16 16 16 16 16 16 -
   Rollout Temperature 1.0 1.0 1.0 1.0 1.0 1.0 -
   Rollout Top-P 1.0 1.0 1.0 1.0 1.0 1.0 -
   Rollout Top-K 50 50 50 50 50 50 -
LLM-as-a-Judge Settings
   Model gpt-4.1-mini gpt-4.1-mini - - - gpt-4.1-mini -
   Temperature 0.4 0.4 - - - 0.4 -

### v2

@@
-Omni-Think: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards
+Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards
 Abstract
@@
 However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning.
-In this work, we introduce Omni-Think, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations.
+In this work, we introduce Omni-Thinker, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations.
 Our approach enables consistent optimization across task types and scales RL-based training to subjective domains.
@@
 1 Introduction
-Figure 1: OMNI-THINK Framework Overview.
+Figure 1: OMNI-THINKER Framework Overview.
 A unified multi-task RL framework for training a single LLM across structured and open-ended tasks.
@@
 Moreover, training LLMs across multiple tasks remains challenging because it requires optimizing for diverse forms of feedback signals, including binary correctness checks in structured tasks and subjective, preference-based judgments in generative ones.
-To address this, we propose Omni-Think, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks.
+To address this, we propose Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks.
 Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks.
@@
 Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
-Our key contributions are: (1) We propose Omni-Think, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains.
+Our key contributions are: (1) We propose Omni-Thinker, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains.
 (2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains.
@@
 4 Methodology
-4.1 Omni-Think Framework Overview
-We introduce Omni-Think, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
+4.1 Omni-Thinker Framework Overview
+We introduce Omni-Thinker, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
 Verifiable Supervision.
@@
 Short-Form Open-Ended Supervision.
-For concise generative tasks with known answers (e.g., General question answering (QA)), we transform the questions into open-ended prompts and embed the distractor responses, which are plausible but incorrect candidate answers, within the context.
-Rather than labeling options, we instruct the model to reason with <think></think> and produce answers in <answer></answer> format.
-A response is assigned a binary reward based on whether it matches any reference answer via string comparison or set membership, promoting semantic grounding and discouraging pattern memorization.
+For language tasks with known or extractable ground-truth answers such as general question answering (QA), we reformulate queries into open-ended prompts and incorporate distractor responses (LLM-generated plausible but incorrect answers) into the context.
+Instead of labeling options, we prompt the model to reason using the <think></think> format and to output answers within <answer></answer> tags.
+Responses are evaluated with a binary reward based on string matching or set membership against reference answers, thereby encouraging semantic grounding and mitigating shallow pattern memorization.
+We find that conditioning the LLM on a diverse set of candidate options, including one correct answer and multiple distractors, is key to steadily improving general-domain reasoning while reducing susceptibility to random guessing or reward hacking, compared to directly prompting the model to generate open-ended answers during training without the augmented context.
 Long-Form Open-Ended Supervision.
@@
 This enables learning in domains where symbolic correctness is insufficient or intractable.
+This prompt-based approach leverages recent advances in the general reasoning capabilities of LLMs, using generated chain-of-thoughts to elicit a ternary reward signal, preferred, tie, or dispreferred, without requiring large-scale preference data collection and reward model training.
 Together, these strategies enable consistent RL training across structured and generative tasks within a unified optimization framework.
@@
 Each entry includes a code prompt and JSON-formatted unit tests for automatic validation.
-For General QA, we extract 5,500 examples from SuperGPQA (M-A-P, 2025), downsampled proportionally by question category.
+For General QA, inspired by SuperGPQA (M-A-P, 2025), we subsample 5,500 queries from the benchmark dataset, proportionally by question category.
 Each sample comprises a factual question paired with a plain-text answer.
-The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on one-turn completions.
+We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM.
+The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions.
 Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
@@
 6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
-We evaluate Omni-Think across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization.
+We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization.
 Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results.
@@
 In Creative Writing, the introduction of our Long-Form Open-Ended Supervision strategy, employing the LLM-as-a-Judge framework for pairwise evaluation, results in significant performance boosts (Curriculum-Guided at 84.2% and Joint MT at 83.00%), underscoring the advantage of our generative reward approach in subjective, open-ended tasks.
-These results support our central hypothesis: The Omni-Think Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
+These results support our central hypothesis: The Omni-Thinker Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
 Figure 3: Average backward transfer (BWT) per target task, showing how training single-task GRPO on all other tasks affects it on average.
@@
 7 Conclusion
-We presented OMNI-THINK, a unified framework that extends LLMs to handle both structured and open-ended tasks within a single policy.
-By integrating reinforcement learning with both rule-based verifiable rewards and generative preference-based supervision, our method improves generalization while addressing challenges such as forgetting and task interference.
+We presented Omni-Thinker, a unified reinforcement learning framework that enables large language models to handle both structured and open-ended tasks under a single policy.
+By combining rule-based verifiable rewards and generative preference-based supervision, our method improves generalization while mitigating forgetting and interference.
 Our findings show that effective multi-task LLM post-training depends not only on reward design but also on how tasks are sequenced and optimized together.
-By ordering tasks from structured to open-ended domains based on backward transfer, OMNI-THINK minimizes forgetting and achieves consistent gains across diverse reasoning and generative tasks.
-Overall, it takes a step toward more general-purpose LLMs that can learn effectively from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation within a single, unified post-training framework.
+Ordering tasks from structured to open-ended domains based on backward transfer reduces forgetting and enhances cross-domain performance.
+Overall, Omni-Thinker advances the goal of general-purpose LLMs that can learn from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation in a single post-training framework.
 Appendix A Appendix
 A.1 Reward Estimation
-Omni-Think employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework.
+Omni-Thinker employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework.
 We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison.

### v3

@@
-Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards
+Omni-Thinker: Scaling Multi-Task RL in LLMs with Hybrid Reward and Task Scheduling
 Abstract
-The advancement of general-purpose artificial intelligence relies on large language models (LLMs) that excel across a wide range of tasks, from structured reasoning to creative generation.
-However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning.
-In this work, we introduce Omni-Thinker, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations.
-Our approach enables consistent optimization across task types and scales RL-based training to subjective domains.
-We further investigate training strategies, demonstrating that a curriculum-based progression that orders tasks from structured to open-ended improves performance and reduces forgetting.
-Experimental results across four domains reveal that curriculum learning improves performance by 5.2% over joint training and 9.1% over model merging.
-These results highlight the importance of task-aware sampling and hybrid supervision in scaling RL-based post-training for general-purpose LLMs.
+The pursuit of general-purpose artificial intelligence depends on large language models (LLMs) that can handle both structured reasoning and open-ended generation.
+We present Omni-Thinker, a unified reinforcement learning (RL) framework that scales LLMs across diverse tasks by combining hybrid rewards with backward-transfer–guided scheduling.
+Hybrid rewards integrate rule-based verifiable signals with preference-based evaluations from an LLM-as-a-Judge, enabling learning in both deterministic and subjective domains.
+Our scheduler orders tasks according to accuracy backward transfer (BWT), reducing forgetting and improving multi-task performance.
+Experiments across four domains show gains of $6.2\%$ over joint training and $12.4\%$ over model merging.
+Moreover, we demonstrate that simple assumptions on accuracy transfer yield accurate predictions of curriculum outcomes, with entropy dynamics explaining deviations due to generative tasks.
+These findings underscore the importance of BWT-aware scheduling and hybrid supervision for scaling RL-based post-training toward general-purpose LLMs.
 1 Introduction
-Figure 1: OMNI-THINKER Framework Overview.
-A unified multi-task RL framework for training a single LLM across structured and open-ended tasks.
-(1) Backward Transfer quantifies task forgettability and informs (2) a curriculum that schedules tasks from least to most forgettable.
-(3) Sampled tasks construct prompts.
-(4) The LLM generates completions, which are (5) scored by hybrid rewards: verifiable (e.g., test case execution), generative (e.g., LLM-as-a-Judge), and auxiliary (e.g., format adherence), among others.
-(6) The policy is updated via Multi-Task GRPO.
-The process repeats via iterative updates (bottom loop).
-As Large Language Models (LLMs) (Hurst et al., 2024; Liu et al., 2024; Dubey et al., 2024; Yang et al., 2024) evolve into general-purpose agents with applications ranging from creative writing (Marco et al., 2025) to robotics (Gemini-Robotics-Team et al., 2025), the challenge of effective post-training becomes increasingly critical.
-While Supervised Fine-Tuning (SFT) is the dominant paradigm for adapting LLMs to downstream tasks, it often encourages memorization rather than robust generalization, especially on tasks that differ from the training distribution or require flexible, context-dependent reasoning (Chu et al., 2025).
-Reinforcement Learning (RL) has emerged as a promising path to improve generalization, especially in structured tasks like math and coding, where verifiable, rule-based rewards can be effective (DeepSeek-AI et al., 2025; Luo et al., 2025; Kimi-Team et al., 2025).
-Methods such as Group Relative Policy Optimization (GRPO) have shown that even coarse signals can guide LLMs toward structured, chain-of-thought responses (Shao et al., 2024).
-However, most RL methods are tailored to deterministic, easily verifiable tasks, limiting their utility in open-ended settings such as question answering and creative writing.
+Figure 1: Performance across four task domains, comparing Joint Training and Curriculum Learning against baselines including SFT and Model Merging.
+Curriculum Learning achieves the strongest results, showing that controlling how tasks are scheduled is crucial for effective multi-task learning.
+Reinforcement learning (RL) has become an effective approach for improving large language models (LLMs) (Hurst et al., 2024; Liu et al., 2024; Dubey et al., 2024; Yang et al., 2024), particularly in structured domains such as math and coding where verifiable, rule-based rewards are available (Guo et al., 2025; Luo et al., 2025; Kimi-Team et al., 2025).
+Methods such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) show that even coarse learning signals can steer LLMs toward structured, chain-of-thought reasoning.
+However, most RL methods remain tailored to deterministically verifiable tasks, limiting their utility in open-ended domains such as general QA and creative writing.
 Moreover, training LLMs across multiple tasks remains challenging because it requires optimizing for diverse forms of feedback signals, including binary correctness checks in structured tasks and subjective, preference-based judgments in generative ones.
-To address this, we propose Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks.
+We address this challenge with Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision under a single policy.
 Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks.
-We further show that curriculum training, which sequences tasks from structured to open-ended based on Backward Transfer (BWT) measurements of task forgettability, significantly improves generalization and reduces forgetting.
-Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
-Our key contributions are: (1) We propose Omni-Thinker, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains.
-(2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains.
-(3) We demonstrate that not all multi-task training strategies are equally effective.
-Curriculum-based task scheduling outperforms uniform sampling and improves generalization in multi-task RL.
-2 Related Work
+Our curriculum is forgetting-aware; it is guided by backward transfer (BWT), where BWT denotes test-performance backward transfer computed on a normalized, task-specific test metric.
+Ordering task training according to this signal yields effective curricula across heterogeneous domains.
+We show that the final accuracy of model after curriculum learning is well predicted by forgettability ranking, even under simplifying assumptions.
+Empirically, we observe complementary entropy dynamics, fine-tuning on creative writing tends to increase the model’s output entropy, whereas training on verifier-supervised, structured tasks tends to decrease it; this trend is consistent with our BWT-guided choice to train structured tasks before open-ended ones.
+Across four domains, Omni-Thinker improves generalization while reducing forgetting, with average gains of 6.2% over joint multi-task training and 12.4% over model merging, respectively.
+Our key contributions are threefold.
+(1) We present Omni-Thinker, a unified framework that trains a single policy across four diverse domains, using hybrid verifiable and preference-based rewards.
+(2) We develop a forgetting-aware curriculum based on backward transfer (BWT) linear ordering maximization over task-specific test performance to reduce forgetting, outperforming joint multi-task training and model merging.
+(3) We empirically analyze training dynamics through the lens of entropy, revealing that structured domains (math, coding) systematically decrease output entropy while open-ended domains (creative writing) increase it, thereby providing an explanatory link between entropy evolution and the effectiveness of BWT-guided curricula.
+2 Framework Overview
+We introduce Omni-Thinker as a unified reinforcement learning framework for large language models that integrates hybrid rewards with task scheduling guided by backward transfer.
+Unlike prior approaches that separate reasoning and generative domains, Omni-Thinker maintains a single policy across heterogeneous tasks, including Math, Coding, General QA, and Creative Writing, while dynamically ordering training to minimize forgetting.
+The framework is instantiated using Multi-Task GRPO, augmented with both symbolic verifiers and LLM-as-a-Judge supervision, and a curriculum determined by accuracy- and entropy-based backward transfers.
+2.1 Notation and Training Objective
+We give ourselves a vocabulary $\mathcal{V}$ with a special end-of-sequence token $\mathrm{eos}$.
+The set of finite sequences of tokens is denoted $\mathcal{V}^{*}$; for any sequence $o\in\mathcal{V}^{*}$, its length is denoted $|o|$ and we say that $o$ is complete if $o_{|o|}=\mathrm{eos}$.
+A model, parameterized by $\theta$, defines a conditional distribution $\pi_{\theta}(y_{t}\mid y_{<t})$ for any given sequence of tokens $(y_{t})_{t\in\mathbb{N}}$.
+It induces a policy $\pi_{\theta}^{\otimes}$ on token sequences defined by $\pi_{\theta}^{\otimes}(o\mid q,o_{<t_{0}}):=\prod_{t=t_{0}}^{|o|}\pi_{\theta}(o_{t}\mid q,o_{<t})$.
+We adopt a multi-task RL (MTRL) formulation: a task is a couple $T=(\mathcal{D},R)$ where $\mathcal{D}$ is a dataset of prompts and $R(q,o)$ is a task-specific reward function.
+Given a set of $K$ tasks $\mathcal{T}=\{T_{1},\ldots,T_{K}\}$, the goal is to learn a unified policy $\pi_{\theta}$ that maximizes the expected reward over the task distribution:
+$\max_{\theta}\;\mathcal{J}(\theta)=\mathbb{E}_{(\mathcal{D},R)\sim P(\mathcal{T})}\left[\mathbb{E}_{q\sim\mathcal{D},o\sim\pi_{\theta}^{\otimes}(\cdot\mid q)}\left[R(o)\right]\right],$ (1)
+where $P(\mathcal{T})$ is a task sampling distribution, which determines task exposure during training.
+In order to train $\pi_{\theta}$ to maximize the objective $\mathcal{J}$, we extend the GRPO (Guo et al., 2025) algorithm to the multi-task setting by jointly optimizing over task-specific reward signals and reference policies.
+For each input prompt $q$, GRPO samples a group of outputs $\{o_{q,1},o_{q,2},\cdots,o_{q,G}\}$ from the old policy $\pi_{\theta_{old}}$.
+A task-specific reward function $R_{k}(q,o)$ scores each output.
+The policy $\pi_{\theta}$ is updated to maximize expected return while controlling divergence from a reference policy.
+We define the policy ratio $\rho_{q,i,t}$ and the normalized advantage estimate $\hat{A}_{q,i,t}$ as follows:
+$\mu_{q}=\text{mean}\big(\{R_{k}(q,o_{q,i})\}_{i=1}^{G}\big),$ $\sigma_{q}=\text{std}\big(\{R_{k}(q,o_{q,i})\}_{i=1}^{G}\big),$ (2)
+$\rho_{q,i,t}=\frac{\pi_{\theta}(o_{q,i,t}\mid q,o_{q,i,<t})}{\pi_{\theta_{\text{old}}}(o_{q,i,t}\mid q,o_{q,i,<t})},$ $\hat{A}_{q,i,t}=\frac{R_{k}(q,o_{q,i})-\mu_{q}}{\sigma_{q}}.$ (3)
+This allows us to write the MT-GRPO objective as
+$\begin{split}&\mathcal{J}_{\text{MT-GRPO}}(\theta)=\mathbb{E}_{k\sim K,q\sim\mathcal{D}_{k},\{o_{q,i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\otimes}(\cdot\mid q_{k})}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{q,i}|}\sum_{t=1}^{|o_{q,i}|}\left\{\min\left[\rho_{q,i,t}\hat{A}_{q,i},\text{clip}\left(\rho_{q,i,t},1{-}\epsilon,1{+}\epsilon\right)\hat{A}_{q,i}\right]-\beta_{k}\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (4)
+where
+$\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]=\frac{\pi_{ref}(o_{q,i,t}|q,o_{q,i,<t})}{\pi_{\theta}(o_{q,i,t}|q,o_{q,i,<t})}-\log\frac{\pi_{ref}(o_{q,i,t}|q,o_{q,i,<t})}{\pi_{\theta}(o_{q,i,t}|q,o_{q,i,<t})}-1.$ (5)
+The clipping parameter $\epsilon$ stabilizes updates by keeping policy ratios within a bounded range, following the PPO approach (Schulman et al., 2017).
+The KL divergence term regularizes the new policy towards the reference policy $\pi_{\text{ref}}$, weighted by a task-specific coefficient $\beta_{k}$.
+2.2 Hybrid Rewards
+We design a hybrid reward system that unifies reinforcement learning across both structured reasoning tasks and open-ended generative domains.
+Verifiable Supervision.
+For tasks with objective correctness signals, such as symbolic math and code generation, we define binary rewards based on symbolic matches, test case results, or other deterministic evaluators depending on the tasks.
+Short-Form Open-Ended Supervision.
+For language tasks with known or extractable ground-truth answers such as general question answering (QA), we reformulate queries into open-ended prompts and incorporate distractor responses (LLM-generated plausible but incorrect answers) into the context.
+Instead of labeling options, we prompt the model to reason using the <think></think> format and to output answers within <answer></answer> tags.
+Responses are evaluated with a binary reward based on string matching or set membership against reference answers, thereby encouraging semantic grounding and mitigating shallow pattern memorization.
+We find that conditioning the LLM on a diverse set of candidate options, including one correct answer and multiple distractors, is key to steadily improving general-domain reasoning while reducing susceptibility to random guessing or reward hacking, compared to directly prompting the model to generate open-ended answers during training without the augmented context.
+Long-Form Open-Ended Supervision.
+For subjective tasks lacking ground truth (e.g., dialogue, writing), we use an LLM-as-a-Judge  (Chen et al., 2025) to assign a scalar reward based on rubric-aligned pairwise preferences between candidate outputs.
+This enables learning in domains where symbolic correctness is insufficient or intractable.
+This prompt-based approach leverages recent advances in the general reasoning capabilities of LLMs, using generated chain-of-thoughts to elicit a ternary reward signal, preferred, tie, or dispreferred, without requiring large-scale preference data collection and reward model training.
+Together, these components form a unified hybrid reward scheme: verifiable rewards ensure correctness where possible and generative-based signals cover subjective domains.
+This design enables reinforcement learning to scale across diverse tasks, from reasoning to open-ended generation.
+2.3 Joint training and Curriculum Learning
+In practice, a maximization step of the training objective $\mathcal{J}_{\text{MT-GRPO}}$ requires a batch $B$ of prompts sampled from $\bigcup_{k=1}^{K}\mathcal{D}_{k}$ then sampling a batch of outputs $\{o_{q,i}\}_{i=1}^{G}$ for each $q\in B$.
+A multi-task schedule is defined as a sequence of batches $(B_{s})_{s=1}^{s_{\max}}$ such that $\forall s\neq s^{\prime},B_{s}\cap B_{s^{\prime}}=\emptyset$ and $\bigcup_{s=1}^{s_{\mathrm{max}}}B_{s}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$.
+Two special cases are considered: Joint Training and Curriculum Learning.
+Joint Training consists in sampling each batch $B$ uniformly at random among all samples (without replacement), disregarding their corresponding tasks: $\forall s,B_{s}\sim\mathcal{U}\left(\bigcup_{k=1}^{K}\mathcal{D}_{k}\setminus\bigcup_{s^{\prime}<s}B_{s^{\prime}}\right)$.
+Curriculum Learning on the other hand consists of pure batches chosen from the same task until exhaustion of the task dataset.
+By pure, we mean that each batch is derived from only one task dataset: $\forall s,B_{s}\sim\mathcal{U}\left(\mathcal{D}_{k_{s}}\setminus\bigcup_{s^{\prime}<s}B_{s^{\prime}}\right)$ for some task schedule $(k_{s})_{s\in\{1,\cdots,s_{\mathrm{max}}\}}$.
+A Curriculum is described by a permutation $\sigma\in\mathfrak{S}_{K}$ of the tasks, with $\mathfrak{S}_{K}$ the set of permutation of $\{1,\cdots,K\}$.
+3 Backward Transfer for Task-scheduling
+We intend to use Backward Transfers (BWT) to guide our choice of curriculum. Following Lopez-Paz & Ranzato (2017) it is defined as follows.
+Definition 1 (Backward Transfer Matrix).
+Let $\theta_{0}$ be a set of initial parameters of a model $\pi_{\theta}$ and let $\theta(\theta_{0},T)$ be the set of parameters obtained after training $\pi_{\theta}$ on task $T$ starting from $\theta_{0}$.
+Write $\mathrm{Acc}(\theta,T)$ the accuracy of model $\pi_{\theta}$ on task $T$.
+The backward transfer matrix is defined by
+$\mathrm{BWT}_{ij}(\theta_{0}):=\log\mathrm{Acc}({\theta(\theta_{0},T_{j})},T_{i})-\log\mathrm{Acc}({\theta_{0}},T_{i}).$ (6)
+Input: $\mathrm{BWT}\in\mathbb{R}^{K\times K}$. $a_{\text{init}}:=\left(\mathrm{Acc}(\theta_{0},T_{k})\right)_{k=1}^{K}\in\mathbb{R}^{K}$. Curriculum $\sigma\in\mathfrak{S}_{K}$
+$a\leftarrow a_{\text{init}}$;
+for $j=1$ to $K$ do
+ $k\leftarrow\sigma(j)$;
+ $a_{k}\leftarrow a_{\text{init},k}$;
+ for $i=1$ to $K$ do
+    $a_{i}\leftarrow a_{i}\times\exp(\mathrm{BWT}_{ik})$;
+   end for
+end for
+return $a$;
+Algorithm 1 Final Accuracy under Assumptions 1 and 2
+3.1 A priori prediction of terminal accuracies under constant BWT
+Our goal is to choose a curriculum order $\sigma$ a priori by predicting the terminal per–task accuracies without training all permutations.
+We propose a simple predictive model in which (i) inter–task backward transfers are treated as constant in log–accuracy, and (ii) training on the full dataset of a task saturates its self–accuracy.
+Under these assumptions, terminal accuracies for any order $\sigma$ become computable from quantities measured once at initialization.
+Setup.
+Using notations from Section 2.3, let $\theta_{0}$ denote the parameters of the pre–trained model $\pi_{\theta}$, and let $\theta_{s}$ be the parameters after $s$ optimization steps following a curriculum order $\sigma\in\mathfrak{S}_{K}$.
+Assumption 1 (Constant off–diagonal BWT in log–accuracy).
+For all $i\neq j$ and all optimization steps $s$ along the schedule,
+$\mathrm{BWT}_{ij}\!\left(\theta_{s}\right)=\mathrm{BWT}_{ij}\!\left(\theta_{0}\right).$ (7)
+Assumption 2 (Task–wise saturation).
+Let $\{B_{s}\}$ be the sequence of mini–batches processed along the schedule.
+If, between steps $s_{1}$ and $s_{2}$, the full dataset $\mathcal{D}_{k}$ of task $T_{k}$ has been seen, then accuracy saturates on task $T_{k}$ to the same accuracy as training from $\theta_{0}$:
+$\bigcup_{s=s_{1}}^{s_{2}}B_{s}=\mathcal{D}_{k}\quad\Rightarrow\quad\mathrm{Acc}\!\left(\theta_{s_{2}},T_{k}\right)=\mathrm{Acc}\!\big(\theta(\theta_{0},T_{k}),\,T_{k}\big).$ (8)
+Theorem 1.
+Under Assumptions 1 and 2, for any curriculum order $\sigma\in\mathfrak{S}_{K}$ starting from $\theta_{0}$, the terminal accuracies $\{\mathrm{Acc}(\theta_{s_{\max}},T_{k})\}_{k=1}^{K}$ given the initialization accuracies $\{\mathrm{Acc}(\theta_{0},T_{k})\}_{k=1}^{K}$, the $\mathrm{BWT}(\theta_{0})$ and a curriculum $\sigma$ is exactly the output of Algorithm 1
+Reasonableness and limitations.
+Assumption 1 abstracts away known drivers of transfer, such as domain overlap, stochasticity, and entropy evolution, thus is a toy yet useful approximation for a priori curriculum selection.
+See Sections 5.2 and 5.3 for a discussion of a correction coming from entropy.
+Working in log–accuracy space keeps accuracies positive but does not eliminate the risk of unrealistic growth over long curricula (e.g., predictions exceeding $1$ in accuracy).
+Assumption 2 is reasonable when each task dataset is sufficiently large and optimization keeps the model in a well–conditioned regime, conditions we satisfy in our experiments.
+Together, these assumptions yield a tractable predictor that captures coarse curriculum effects while remaining simple enough to evaluate without exhaustive training.
+3.2 Curriculum Choice via Linear Ordering Maximization
+Algorithm 1 admits the following closed form for the predicted terminal accuracies:
+$\log\mathrm{Acc}(\sigma)-\log\mathrm{Acc}(\mathrm{Id})=\sum_{i<j}\bigl(\Phi_{\sigma}^{-1}\,\mathrm{BWT}\,\Phi_{\sigma}\bigr)_{ij},\quad\text{with}\quad\Phi_{\sigma,ij}=\mathbf{1}_{i=\sigma(j)}.$ (9)
+Input: BWT matrix $\mathrm{BWT}\in\mathbb{R}^{K\times K}$
+$\sigma\leftarrow$ empty list;
+while there are unvisited tasks do
+ $k^{*}\leftarrow\arg\max_{k\notin\sigma}\ \sum_{i\notin\sigma\cup\{k\}}\mathrm{BWT}_{ik}$;
+   append $k^{*}$ to $\sigma$;
+end while
+return $\sigma$;
+Algorithm 2 Greedy BWT-LOM Curriculum
+In words, curriculum reordering acts by permuting the BWT matrix with $\Phi_{\sigma}$, and the gain relative to the identity schedule is simply the sum of the upper–triangular entries of the permuted matrix.
+Given an aggregated score of the form
+$\mathcal{S}:=\sum_{T\in\mathcal{T}}\alpha_{T}\log\mathrm{Acc}(\theta,T),$ (10)
+identifying the best task order amounts to solving a Linear Ordering Problem (LOP).
+see (Floudas & Pardalos, 2008) for an overview.
+This problem is known to be NP–hard, but for small numbers of tasks ($K$) it can be solved exactly.
+For larger $K$, a wide range of approximation algorithms and heuristics exist.
+A simple heuristic is to rank tasks by a forgettability score: $F_{k}:=\alpha_{k}\sum_{i\neq k}\mathrm{BWT}_{ik}$.
+Intuitively, ordering tasks by decreasing $F_{k}$ prioritizes those that exert the least destructive interference on others (or even provide positive transfer), thereby reducing overall forgetting.
+Our curriculum orders tasks by decreasing column mean of BWT.
+4 Experimental Setup
+Training Datasets.
+We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation.
+For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning.
+We further subsample 12,000 examples to fit our compute budget.
+For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed.
+Each entry includes a code prompt and JSON-formatted unit tests for automatic validation.
+For General QA, we subsample 5,500 queries from from SuperGPQA (Du et al., 2025) dataset, proportionally by question category.
+Each sample comprises a factual question paired with a plain-text answer.
+We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM.
+The Creative Writing.
+domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions.
+Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
+Evaluation.
+We assess performance in each domain using dedicated benchmarks aligned with the task’s evaluation criteria.
+Math: accuracy on AIME24 (MAA, 2024), AMC23 (MAA, 2023), Gaokao2023EN (Liao et al., 2024), MATH-500 (Hendrycks et al., 2021), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024).
+Coding: pass@1 on BigCodeBench (Complete-Full) (Zhuo et al., 2024) and LiveCodeBench (24Oct–25Jan) (Jain et al., 2024).
+General QA: exact-match accuracy on MMLU-Pro (Wang et al., 2024).
+Creative Writing: win rate on the role-play and creative writing subsets of MT-Bench (Zheng et al., 2023), against GPT-4 (pre-gen, June 16, 2023).
+Baselines.
+We use Qwen2.5-7B-Instruct (Yang et al., 2024) as the base model for all experiments, owing to its strong instruction-following capability, which makes it well-suited for reinforcement learning on both structured reasoning and open-domain QA tasks.
+Supervised Fine-Tuning (SFT): In order to have a meaningful comparison with GRPO, we adopt a similar self-sampled data curation and fine-tuning approach with Rejection sampling Fine-Tuning (Yuan et al., 2023).
+We first prompt the base model to generate 128 chain-of-thought responses for our training dataset to ensure we end up with at least one correct response for most queries, then filter them based on the same accuracy reward signals used in GRPO training.
+We then perform sft on base model using these self-distilled responses.
+This provides a strong on-policy learning baseline that incorporates explicit reasoning steps through self-distillation from the base model.
+Model Merging: We employ TIES-Merging  (Wu et al., 2025) as our model‑merging baseline.
+It is a simple yet effective method designed specifically for the multi-task setting that takes into consideration the interference between parameters from models trained on individual tasks during the merging process.
+It has demonstrated superior performance in multi-task learning compared to linear and task arithmetic approaches (Yadav et al., 2023).
+To begin with, we conduct single-task GRPO training using individual task datasets and collect the model weights of the best checkpoints with the help of a validation set for each training run.
+We then merge the four single-task models using a scaling value $\lambda=1$.
+5 Results and Discussion
+5.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
+We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization.
+BWT matrix is computed following equation 6, then Algorithm 1 is used to predict the accuracy of the model after curriculum learning, Appendix A.2.2 for details.
+The predicted best curriculum using Algorithm 2 is Code $\rightarrow$ Math $\rightarrow$ QA $\rightarrow$ Writing while the worst is Writing $\rightarrow$ QA $\rightarrow$ Math $\rightarrow$ Coding.
+Figure 1 shows that Curriculum Multi-Task Learning with GRPO consistently yields the best results. Table 1 further details how these gains vary by benchmarks.
+In Math, Curriculum Learning (CL) achieves the highest average performance at 59.6%, with the clearest gains on more complex reasoning tasks such as MinervaMath and OlympiadBench.
+These benchmarks benefit from strong rule-based reward signals and backward-transfer-guided task ordering.
+In contrast, datasets like AMC23 show minimal change because their relatively high baseline scores likely reflect smaller question sets and potential pretraining overlap rather than robust multi-step problem-solving.
+In General QA, CL again performs best (52.2%), followed by Model Merging (49.8%) and Mixed Training GRPO (48.8%).
+These improvements are driven by our Short-Form Open-Ended Supervision strategy: instead of generating responses in a fully open-ended and unconstrained fashion, the model is trained to produce complete answer strings given a diverse set of candidate responses, enabling the effective application of verifiable reward through simple string matching when training general-domain tasks.
+For Code Generation, CL achieves 35.4%, slightly ahead of Model Merging.
+Notably, we only evaluate on the subset of LiveCodeBench(24Oct-25Jan) problems released after Qwen2.5’s data cutoff, which ensures that these are unseen test items.
+This setup highlights CL’s significant generalization gains on novel problems, explaining the larger improvements on LiveCodeBench relative to static benchmarks like BigCodeBench, where data overlap is more likely.
+In Creative Writing, the introduction of our Long-Form Open-Ended Supervision strategy, employing the LLM-as-a-Judge framework for pairwise evaluation, results in significant performance boosts (Curriculum-Guided at 84.2% and Joint MT at 83.00%), underscoring the advantage of our generative reward approach in subjective, open-ended tasks.
+These results support our central hypothesis: The Omni-Thinker Framework, BWT-guided Curriculum Learning with hybrid rewards, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
+Table 1: Performance across benchmarks.
+ST = Single-Task RL (e.g., ST Math = RL trained only on math).
+$M$ = Model Merging.
+JT = Joint Training.
+CL = Curriculum Learning.
+Bolded values mark the best per row; underscored values mark the second best.
+Domains include Math (7 sets), MMLU-Pro (9 categories), Coding (2 sets), and Creative Writing (MT-Bench).
+[…]
+Math
+AIME24 18.0 13.3 14.7 14.0 15.3 16.7 10.0 11.3 15.3
+AMC23 57.5 57.5 60.0 62.5 61.0 62.0 56.0 51.0 70.0
+Gaokao2023en 73.0 74.3 76.1 74.0 75.6 74.3 74.8 76.6 77.1
+MATH500 78.2 78.8 80.4 75.4 79.2 76.8 79.8 77.6 81.0
+MinervaMath 64.3 64.0 66.5 63.2 61.8 65.1 66.2 68.4 71.7
+OlympiadBench 42.1 43.0 43.7 41.3 43.0 43.0 41.8 43.6 47.4
+Average 55.5 55.1 56.9 55.1 56.0 56.3 54.8 54.8 60.4
+General QA
+Biology 57.6 56.8 52.3 67.4 59.0 66.3 65.6 67.2 68.8
+Business 33.5 39.0 25.6 58.7 33.0 48.2 59.8 49.8 47.5
+Chemistry 35.8 31.8 27.3 47.7 38.3 44.1 42.5 42.1 50.7
+CS 53.7 48.1 50.2 55.1 52.0 53.7 53.9 58.8 59.3
+Economics 42.7 49.2 38.7 62.9 44.9 59.6 62.0 56.8 62.1
+Engineering 28.3 31.3 20.4 37.5 26.6 37.8 38.1 35.8 37.1
+Health 46.7 46.2 45.2 51.0 47.2 45.7 52.7 50.7 57.1
+History 37.3 33.3 34.7 47.2 38.6 33.9 47.3 43.3 45.7
+Law 23.2 24.0 20.6 27.9 23.3 26.8 26.6 27.5 29.7
+Math 55.4 52.6 50.4 59.3 56.3 57.4 58.3 59.2 61.2
+Other 44.3 40.0 39.7 51.0 43.9 46.4 51.8 49.9 53.3
+Philosophy 36.9 34.3 33.3 43.9 35.5 38.2 41.5 42.1 42.9
+Physics 41.1 37.4 30.8 53.7 41.6 49.8 46.7 48.0 55.6
+Psychology 50.9 51.5 45.4 60.2 51.8 59.0 59.4 59.3 61.8
+Average 41.5 40.1 37.9 51.3 42.0 47.8 49.8 48.8 52.2
+Coding
+BigCodeBench 46.5 50.4 46.7 47.1 46.8 44.5 48.1 47.2 49.5
+LiveCodeBench 12.7 21.8 13.1 13.8 13.3 14.2 20.8 17.0 21.3
+Average 29.6 36.1 29.9 30.5 30.1 29.3 34.4 32.1 35.4
+Creative Writing
+MT-Bench (Writing) 74.2 71.6 74.2 63.0 78.3 44.2 67.5 83.0 84.2
+5.2 Entropy Dynamics: Discussion
+Comparison between accuracies predictions to the actual test evaluation results for various curricula is depicted on Table 3.
+Predicted accuracies using test set backward transfers are surprisingly precise considering our simplifying assumptions, especially for the top curriculum.
+We now discuss an identified cause of discrepancies.
+We define the token-wise entropy of a policy $\pi_{\theta}$ on task $T_{k}$ ti measures the average per-token uncertainty of the policy across task samples.
+$\mathcal{E}(\theta,T_{k}):=-\mathbb{E}_{q\sim\mathcal{D}_{k},o\sim\pi_{\theta}^{\otimes}(\cdot\mid q)}\frac{1}{|o|}\sum_{t=1}^{|o|}\sum_{v\in\mathcal{V}}\pi_{\theta}(v\mid q,o_{<t})\log\pi_{\theta}(v\mid q,o_{<t}).$ (11)
+Table 2: Test performance (%) of single-task RL fine-tuning on General QA and Creative Writing respectively under different generation temperature (T) in training.
+Eval Task General QA Creative Writing
+T=1.0 T=1.2 T=1.0 T=0.1
+Math 55.1 54.8 55.6 57.8
+Coding 30.4 26.6 30.1 27.4
+General QA 51.4 13.9 42.0 44.9
+Creative Writing 63.0 66.7 78.3 82.5
+Entropy has been shown to drop during GRPO fine-tuning on reasoning, coding and more generally verified rewards (Rastogi et al., 2025; Cui et al., 2025; Yu et al., 2025) with a correlation to higher accuracy until reaching a breaking point.
+Long training requires extra care towards entropy scaling either via regularization (Shen, 2025) or dynamic temperature scaling.
+Our multi-task setting differs in two key aspects compared to the above references.
+First, we are using hybrid rewards with both verified and generative components.
+The Creative Writing task is generative and is expected to increase entropy.
+Indeed, Wang et al.
+(2025) account for so-called forking tokens corresponding to structural choices of the output capturing most of the entropy.
+Reasoning tasks tend to require highly causal token sequences hence few forking tokens (low entropy) while generative tasks may allow more logical cuts at inference (higher entropy).
+Second, we train on multiple domains compared to mostly single-domain analysis in the references above.
+It is unclear a priori whether entropy decrease propagates from task to task.
+Agarwal et al.
+(2025) show that fine-tuning to reduce entropy suffices to improve performance on multiple domains.
+We hypothesize that models implicitly learn to emulate lower or higher temperatures as a mechanism to regulate entropy.
+In practice, the policy often produces logically flawed but rarely syntactically meaningless outputs, suggesting that its support lies within a constrained domain $V(q,o_{<t})\subset\mathcal{V}$: $\sum_{v\in V(q,o_{<t})}\pi_{\theta}(v\mid q,o_{<t})=1$.
+Scaling the final layer weights by a factor $\lambda<1$ preserves this domain while increasing entropy, effectively raising the model’s base temperature.
+Such changes propagate across all tasks, not only the one being fine-tuned.
+Thus, even when two domains are sufficiently distinct for knowledge transfer to fail, the entropy dynamics may still be measurable across tasks.
+Table 3: Comparison of empirical and predicted test accuracies (%). Each task column reports Test vs. Predicted accuracy for a given curriculum order. Standard deviations are rounded up.
+Curriculum Math Coding QA Writing
+Test Pred Test Pred Test Pred Test Pred
+$\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ $60.4\!\pm\!0.3$ $57.3$ $\pm\!0.3 35.435.4$ $52.2\!\pm\!0.1$ $51.9$ $84\!\pm\!2$ $78.4$
+$\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ $9.3\!\pm\!0.3 55$.655.6 $31.6\!\pm\!0.3$ $36.7$ $39.0\!\pm\!0.1$ $38.8$ $79\!\pm\!2$ $78.4$
+$\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ $60.4\!\pm\!0.3$ $57.6$ $1.9\!\pm\!0.3 33$.933.9 $36.3\!\pm\!0.1$ $38.8$ $82\!\pm\!2$ $73.1$
+$\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ $6.6\!\pm\!0.3 55$.555.5 $32.7\!\pm\!0.3$ $36.1$ $22.6\!\pm\!0.1$ $38.4$ $75\!\pm\!2$ $62.1$
+5.3 Entropy Dynamics: Empirical Support
+Figure 2: Validation Set Entropy Change Matrix.
+The intuitions laid out in the previous section are empirically supported by two experiments.
+We define the entropy change matrix as
+$H_{ij}:=\frac{\mathcal{E}(\theta(\theta_{0},T_{i}),T_{j})}{\mathcal{E}(\theta_{0},T_{j})}-1$ (12)
+and compute it, see Figure 2.
+We observe that Math and Coding decrease entropy for all tasks (as previously observed for Verified Rewards) while Creative Writing increases entropy.
+Also, entropy change seems to depend primarily on the source task type, secondarily on the target task type, but not on their domain overlap.
+We fine-tune the base model on QA and Writing task with different choices of temperature to emulate the effect of entropy modifications due to fine-tuning on entropy-increasing or entropy-decreasing tasks.
+On the one hand, Writing is thought to benefits from temperature lowering coming from other tasks, we thus train the base model on writing with a lower temperature and expect the model to close the gap compared to the best Curriculum-trained model when evaluated with zero temperature.
+On the other hand, QA is thought to performs worse than expected in the worst curriculum due to the increased entropy coming from the trainig on Writing.
+We fine-tune the base model on QA with a higher temperature, then evaluate with zero temperature, and expect QA performance to drop toward the low performance of the worse curriculum learning.
+The results on table 2 shows that indeed the case.
+6 Related Work & limitations
 Large Language Models and Multi-Task Learning
@@
 Moreover, our curriculum design, guided by backward transfer, helps maintain stable cross-task performance even for tasks lacking deterministic evaluation criteria.
-3 Problem Formulation
-We consider the problem of training a single language model policy to perform well across a diverse set of tasks, including mathematical reasoning, code generation, question answering, and creative writing.
-Let $\mathcal{T}=\{T_{1},\ldots,T_{K}\}$ be a collection of $K$ tasks, where each task $T_{k}$ has a dataset $\mathcal{D}_{k}$ of input-output pairs $(q_{k},o^{*}_{k})$, where $q_{k}\in\mathcal{Q}_{k}$ is an input prompt, and $o^{*}_{k}\in\mathcal{O}^{*}_{k}$ is the corresponding ground-truth output response.
-The model, parameterized by $\theta$, defines a conditional distribution $\pi_{\theta}(y_{k}\mid x_{k})$.
-Supervised fine-tuning (SFT) aims to maximize the likelihood of ground-truth outputs by minimizing:
-$\min_{\theta}\sum_{k=1}^{K}\mathbb{E}_{(x_{k},y_{k})\sim\mathcal{D}_{k}}\left[\sum_{t=1}^{T}-\log\pi_{\theta}(o^{*}_{k,t}\mid q_{k},o^{*}_{k,<t})\right],$ (1)
-where $o^{*}_{k,t}$ denotes the $t$-th token of $o^{*}_{k}$, and $T$ is the number of tokens in the output.
-While effective in-domain, SFT often overfits to training data and fails to generalize, particularly when task formats vary.
-Recent studies (Chu et al., 2025) show that reinforcement learning (RL) with outcome-based rewards better promotes cross-task generalization.
-Motivated by this, we adopt a multi-task RL (MTRL) formulation with task-specific reward functions $R_{k}$.
-The goal is to learn a unified policy $\pi_{\theta}$ that maximizes the expected reward over the task distribution:
-$\max_{\theta}\;\mathcal{J}(\theta)=\mathbb{E}_{T_{k}\sim P(\mathcal{T})}\left[\mathbb{E}_{q_{k}\sim\mathcal{D}_{k},o_{k}\sim\pi_{\theta}(\cdot\mid x_{k})}\left[R_{k}(o_{k})\right]\right],$ (2)
-where $P(\mathcal{T})$ is the task sampling distribution, which determines task exposure during training, and $R_{k}$ is the task-specific reward function.
-This objective trains a single policy, $\pi_{\theta}$, that balances performance across tasks, enabling knowledge transfer, mitigating negative transfer, and improving generalization across domains.
-4 Methodology
-4.1 Omni-Thinker Framework Overview
-We introduce Omni-Thinker, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
-Verifiable Supervision.
-For tasks with objective correctness signals, such as symbolic math and code generation, we define binary rewards based on symbolic matches, test case results, or other deterministic evaluators depending on the tasks.
-Short-Form Open-Ended Supervision.
-For language tasks with known or extractable ground-truth answers such as general question answering (QA), we reformulate queries into open-ended prompts and incorporate distractor responses (LLM-generated plausible but incorrect answers) into the context.
-Instead of labeling options, we prompt the model to reason using the <think></think> format and to output answers within <answer></answer> tags.
-Responses are evaluated with a binary reward based on string matching or set membership against reference answers, thereby encouraging semantic grounding and mitigating shallow pattern memorization.
-We find that conditioning the LLM on a diverse set of candidate options, including one correct answer and multiple distractors, is key to steadily improving general-domain reasoning while reducing susceptibility to random guessing or reward hacking, compared to directly prompting the model to generate open-ended answers during training without the augmented context.
-Long-Form Open-Ended Supervision.
-For subjective tasks lacking ground truth (e.g., dialogue, writing), we use an LLM-as-a-Judge  (Chen et al., 2025) to assign a scalar reward based on rubric-aligned pairwise preferences between candidate outputs.
-This enables learning in domains where symbolic correctness is insufficient or intractable.
-This prompt-based approach leverages recent advances in the general reasoning capabilities of LLMs, using generated chain-of-thoughts to elicit a ternary reward signal, preferred, tie, or dispreferred, without requiring large-scale preference data collection and reward model training.
-Together, these strategies enable consistent RL training across structured and generative tasks within a unified optimization framework.
-4.2 Multi-Task Group Relative Policy Optimization
-We extend the GRPO (Guo et al., 2025) algorithm to the multi-task setting by jointly optimizing over task-specific reward signals and reference policies.
-For each input prompt $q_{k}$, GRPO samples a group of outputs $\{o_{k,1},o_{k,2},\cdots,o_{k,G}\}$ from the old policy $\pi_{\theta_{old}}$.
-At each generation step $t$, the partial output is denoted $o_{k,i,<t}$, where $o_{k,i,<t}=[o_{k,i,1},\dots,o_{k,i,t-1}]$, that is, the input concatenated with the sequence of generated tokens up to step $t-1$.
-A reward function $R_{k}(o,\phi_{k})$ scores each output,where $\phi_{k}$ denotes task-specific information required for the evaluation, such as reference answers, test cases, or comparison outputs required for evaluation.
-The policy $\pi_{\theta}$ is updated to maximize expected return while controlling divergence from a reference policy.
-Let
-$\mu_{k}=\text{mean}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},\quad\sigma_{k}=\text{std}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},$ (3)
-we define the policy ratio $\rho_{k,i,t}$ and the normalized advantage estimate $\hat{A}_{k,i,t}$ as
-$\rho_{k,i,t}=\frac{\pi_{\theta}(o_{k,i,t}\mid q_{k},o_{k,i,<t})}{\pi_{\theta_{\text{old}}}(o_{k,i,t}\mid q_{k},o_{k,i,<t})},$ $\quad\hat{A}_{k,i,t}=\frac{R_{k}(o_{k,i},\phi_{k})-\mu_{k}}{\sigma_{k}}.$ (4)
-This allows us to write the MT-GRPO objective as
-$\begin{split}&\mathcal{J}_{\text{MT-GRPO}}(\theta)=\mathbb{E}_{k\sim K,q_{k}\sim\mathcal{D}_{k},\{o_{k,i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{k})}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{k,i}|}\sum_{t=1}^{|o_{k,i}|}\left\{\min\left[\rho_{k,i,t}\hat{A}_{k,i},\text{clip}\left(\rho_{k,i,t},1{-}\epsilon,1{+}\epsilon\right)\hat{A}_{k,i}\right]-\beta_{k}\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (5)
-where
-$\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]=\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1.$ (6)
-The clipping parameter $\epsilon$ stabilizes updates by keeping policy ratios within a bounded range, following the PPO approach (Schulman et al., 2017).
-The KL divergence term regularizes the new policy towards the reference policy $\pi_{\text{ref}}$, weighted by a task-specific coefficient $\beta_{k}$.
-4.3 Task Scheduling
-We consider two strategies for optimizing a shared policy across multiple tasks: curriculum learning and joint training.
-Both utilize task-specific rewards but differ in scheduling.
-To guide task ordering, we adapt a Backward Transfer (BWT) metric to quantify forgetting in our multi-task setting.
-Following Lopez-Paz & Ranzato (2017), we define $BWT_{j}=P_{after,j}-P_{base,j}$, where $P_{base,j}$ is the performance on task $j$ from the original base model and $P_{after,j}$ is the performance on task $j$ after training on subsequent tasks.
-A negative $BWT_{j}$ indicates that performance on task $j$ degrades due to learning new tasks, measuring the extent of forgetting.
-Curriculum Learning:
-Prior work in curriculum learning prioritizes task complexity (e.g., easy to hard) for more effective learning (Bengio et al., 2009; Parashar et al., 2025).
-We instead adopt a forgetting-based curriculum that orders tasks according to their average forgettability in multi-task learning.
-Tasks with the least negative or even positive backward transfer are introduced earlier, serving as stable foundations for subsequent tasks that are more prone to catastrophic forgetting.
-In our setup, this coincides with increasing supervision subjectivity.
-This forgetting-based ordering reduces overall forgetting and improves generalization by minimizing destructive interference across tasks.
-Joint Training:
-All tasks are sampled concurrently from a shared distribution.
-While this promotes exposure, it introduces three key challenges: (1) simpler tasks may dominate without adaptive sampling due to the tendency of reward hacking in GRPO; (2) reward delays vary, requiring asynchronous updates; and (3) Kullback–Leibler (KL) regularization demands tuning per task due to varying sensitivities.
-In both regimes, we apply MT-GRPO with normalized advantages and task-specific KL coefficients.
-At each step, a task $T_{k}$ is selected, completions are generated by the current policy $\pi_{\theta}$, and updates are applied based on the corresponding reward signal.
-5 Experimental Setup
-5.1 Training Datasets
-We curate a multi-domain training dataset covering Math, Coding, General QA, and Creative Writing, with each domain selected to support hybrid reward functions and robust evaluation.
-For Math, we begin with the OpenR1-Math (HuggingFace, 2025) dataset, retaining only word problems and excluding questions that require visual reasoning.
-We further subsample 12,000 examples to fit our compute budget.
-For Coding, data is sourced from the code-r1-12k (Liu & Zhang, 2025) dataset, with outliers exceeding 1024 tokens removed.
-Each entry includes a code prompt and JSON-formatted unit tests for automatic validation.
-For General QA, inspired by SuperGPQA (M-A-P, 2025), we subsample 5,500 queries from the benchmark dataset, proportionally by question category.
-Each sample comprises a factual question paired with a plain-text answer.
-We then generate 15 additional confusion options while making sure the uniqueness of correctness by prompting an LLM.
-The Creative Writing domain leverages 6,650 conversations from Nitral AI’s ShareGPT dataset (Nitral‑AI, 2024), focused on single-turn completions.
-Samples exceeding two dialogue turns are filtered out, and responses are judged via an LLM-as-a-Judge framework.
-5.2 Baselines
-We adopt Qwen2.5-7b-Instruct as the base model for all our experiments (Yang et al., 2024).
-Its robust instruction-following ability makes it a suitable candidate for subsequent reinforcement learning training on both reasoning tasks and more general open-domain QA, as it relies on the model’s capacity to comprehend and follow the given prompts effectively.
-Supervised Fine-Tuning (SFT):
-In order to have a meaningful comparison with GRPO, we adopt a similar self-sampled data curation and fine-tuning approach with Rejection sampling Fine-Tuning (Yuan et al., 2023).
-We first prompt the base model to generate 128 chain-of-thought responses for our training dataset to ensure we end up with at least one correct response for most queries, then filter them based on the same accuracy reward signals used in GRPO training.
-We then perform supervised fine-tuning on Qwen2.5-7b-Instruct using these self-distilled responses.
-This provides a strong on-policy learning baseline that incorporates explicit reasoning steps through self-distillation from the base model.
-Model Merging:
-We employ TIES-Merging  (Yadav et al., 2023b) as our model‑merging baseline.
-It’s a simple yet effective method designed specifically for the multi-task setting that takes into consideration the interference between parameters from models trained on individual tasks during the merging process.
-It has demonstrated superior performance in multi-task learning compared to linear and task arithmetic approaches (Yadav et al., 2023a).
-To begin with, we conduct single-task GRPO training using individual task datasets and collect the model weights of the best checkpoints with the help of a validation set for each training run.
-We then merge the four single-task models using a scaling value $\lambda=1$.
-5.3 Evaluation
-We assess performance in each domain using dedicated held-out benchmarks aligned with the task’s unique evaluation criteria. The detailed evaluation set is presented as follows
-Math Reasoning:
-we compute accuracy over seven datasets: AIME24 (MAA, 2024), AMC23 (MAA, 2023), Gaokao2023EN (Liao et al., 2024), MATH-500 (Hendrycks et al., 2021), MinervaMath (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024).
-Code Generation:
-we measure coding ability via pass@1 on BigCodeBench (Complete-Full) (Zhuo et al., 2024) and LiveCodeBench (24Oct-25Jan) (Jain et al., 2024).
-General QA:
-we report exact-match accuracy using the MMLU-Pro benchmark (Wang et al., 2024).
-Creative Writing:
-we evaluate the creative writing task using the role-play and creative writing subcategory of MT‑Bench (Zheng et al., 2023), reporting win rate against a GPT-4 (pre-gen dated June 16, 2023) model.
-6 Results and Discussion
-Figure 2: Performance gains across four task domains, comparing our Multi-Task (MT) framework (Joint and Curriculum variants) against baselines including Supervised Fine-Tuning (SFT) and Model Merging.
-Curriculum MT achieves the strongest results, particularly in open-ended tasks, showing that controlling how multi-task learning is structured is crucial for effective generalization.
-6.1 Main Results: Scaling Multi-Task LLM Post-Training with Omni-Think
-We evaluate Omni-Thinker across four diverse domains: Coding, Math, General QA, and Creative Writing, to assess how reinforcement learning with rule-based verifiable rewards and generative supervision supports multi-task generalization.
-Figure 2 shows that Curriculum-Guided Multi-Task GRPO consistently yields the best results.
-Table 2 further details how these gains vary by benchmarks.
-In Math, Curriculum MT achieves the highest average performance at 59.6%, with the clearest gains on more complex reasoning tasks such as MinervaMath and OlympiadBench.
-These benchmarks benefit from strong rule-based reward signals and backward-transfer-guided task ordering.
-In contrast, datasets like AMC23 show minimal change because their relatively high baseline scores likely reflect smaller question sets and potential pretraining overlap rather than robust multi-step problem-solving.
-In General QA, Curriculum MT again performs best (52.2%), followed by Model Merging (49.8%) and Joint Multi-Task GRPO (48.8%).
-These improvements are driven by our Short-Form Open-Ended Supervision strategy: instead of generating responses in a fully open-ended and unconstrained fashion, the model is trained to produce complete answer strings given a diverse set of candidate responses, enabling the effective application of verifiable reward through simple string matching when training general-domain tasks.
-For Code Generation, Curriculum MT achieves 35.4%, slightly ahead of Model Merging.
-Notably, we only evaluate on the subset of LiveCodeBench(24Oct-25Jan) problems released after Qwen2.5’s data cutoff, which ensures that these are unseen test items.
-This setup highlights Curriculum MT’s significant generalization gains on novel problems, explaining the larger improvements on LiveCodeBench relative to static benchmarks like BigCodeBench, where data overlap is more likely.
-In Creative Writing, the introduction of our Long-Form Open-Ended Supervision strategy, employing the LLM-as-a-Judge framework for pairwise evaluation, results in significant performance boosts (Curriculum-Guided at 84.2% and Joint MT at 83.00%), underscoring the advantage of our generative reward approach in subjective, open-ended tasks.
-These results support our central hypothesis: The Omni-Thinker Training Framework, with the help of Curriculum MT, enables a single unified policy to scale across structured and open-ended tasks alike, without relying on interleaving RLVR on reasoning tasks and fine-tuning non-reasoning tasks.
-Figure 3: Average backward transfer (BWT) per target task, showing how training single-task GRPO on all other tasks affects it on average.
-A negative BWT indicates forgetting; a positive value indicates positive transfer.
-Figure 4: Models trained to generate full-text answers outperform those trained to select letter choices. This format promotes deeper semantic understanding over shallow pattern matching or guessing.
-6.2 Training Order Matters: From Least to Most Forgettable
-Our analysis highlights that task ordering is critical for robust multi-task reinforcement learning.
-Figure 4 summarizes the average backward transfer (BWT) each task receives when training single-task GRPO models on other tasks.
-Specifically, the x-axis of Figure 4 represents the target task.
-A negative average BWT value for a target task indicates that it tends to suffer forgetting or experience a decrease in performance when other tasks are trained before it; a positive value suggests that the target task tends to benefit or experience positive transfer.
-This high-level view reveals that Creative Writing is the most vulnerable task, as it experiences the most negative average BWT received, while Coding is the most resilient, showing a positive average BWT received.
-For example, Coding benefits on average after training on another task (average BWT received $+1.06\%$, indicating a positive transfer).
-In contrast, Creative Writing is the most vulnerable to forgetting: it has the highest magnitude of negative average BWT received ($-4.57\%$), meaning it easily loses performance when other tasks are trained first.
-Figure 5 shows the full BWT matrix, where x-axis consists of target tasks and y-axis represent source tasks that are being trained on.
-Each cell $(row_{i},col_{j})$ in the matrix represents the BWT from source task $j$ to target task $i$.
-A negative value indicates that training on source task $j$ causes forgetting in target task $i$, while a positive value suggests positive transfer.
-We observe that structured tasks like Coding and Math are both less forgetting conducive as sources and tend to receive more positive transfers as targets from training on other tasks.
-Guided by this analysis, our curriculum trains tasks in a least-to-most forgettable order: Coding $\rightarrow$ Math $\rightarrow$ General QA $\rightarrow$ Creative Writing.
-This progression grounds the model in stable, verifiable tasks before introducing those more prone to interference and forgetting.
-As confirmed by Table 1, reversing this order significantly harms performance.
-Overall, our findings suggest that BWT patterns can offer a practical, empirically informed guide for task sequencing, helping reduce negative transfer and improve generalization in multi-task RL for LLMs.
-Further work is needed to test this approach across a broader range of tasks and domains, including those that require logical reasoning over graph-structured data  (Zhou et al., 2024) and knowledge-base retrieval  (Dehghan et al., 2024), as well as more diverse open-ended domains such as mixed-initiative collaborative storytelling and co-creativity  (Kreminski et al., 2024).
-6.3 Output Format Matters: Full-Text Answers Enhance Generalization
-We examine how output format impacts generalization by comparing models trained to generate full-text answers versus selecting letter choices in multiple-choice QA (MCQ).
-Using GRPO, we train two single-task policies on the training set, one prompted to produce full-text final answers at the end of its chain-of-thought completions, and the other to output only letter choices (e.g., “A”, “B”, “C”).
-As shown in Figure 4, the model trained to output full-text answer achieves significantly better generalization when evaluated with free-form QA prompts on MMLU Pro (51% vs.
-41%).
-While the letter-choice model slightly outperforms when evaluated strictly on MCQ prompts, the full-text model remains competitive across both prompt formats.
-These results suggest that training with complete, semantically grounded answers encourages deeper reasoning, improving the model’s ability to generalize beyond the specific format seen during training.
-In contrast, letter-choice training risks overfitting to shallow pattern matching, reducing transferability to realistic QA settings that often require articulated responses.
 7 Conclusion
@@
 Overall, Omni-Thinker advances the goal of general-purpose LLMs that can learn from both verifiable and subjective feedback, bridging structured reasoning, open-ended question answering, and creative generation in a single post-training framework.
+Limitations:
+Further work is needed to test this approach across a broader range of tasks and domains, including those that require logical reasoning over graph-structured data  (Zhou et al., 2024) and knowledge-base retrieval  (Dehghan et al., 2024), as well as more diverse open-ended domains such as mixed-initiative collaborative storytelling and co-creativity  (Kreminski et al., 2024).
+Our discussion on entropy is consistent with the existing literature however we lack 1) experiments backing our intuition that forking tokens are clearly identifiable and more frequent in generative tasks is lacking 2) a direct measurement of temperature scaling due to fine-tuning.
+Furthermore, our discussion on entropy is mostly qualitative.
+A quantitative study of entropy effects could allow to avoid failure cases of our accuracy predictions hence of our task scheduler.
+Finally, we fully leverage BWT only for curriculum learning.
+A comprehensive framework combining BWT, entropy and gradient norm in a quantitative manner in the continuous limit would allow for dynamical joint training with mixed batches.
 Appendix A Appendix
 A.1 Reward Estimation
-Omni-Thinker employs a hybrid reward system combineing rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework.
-We define task-specific reward functions as $r_{k}(o,\phi_{k})$, where $o$ denotes the model output and $\phi_{k}$ represents auxiliary task-specific inputs required for evaluation—such as reference answers, unit tests, or candidate responses for comparison.
+Omni-Thinker employs a hybrid reward system combining rule-based correctness (math, code, QA) with preference-based supervision (creative writing) in a unified RL framework.
+We define task-specific reward functions as $R_{k}(q,o)$, where $o$ denotes the model output and $q$ is the prompt provided to the model.
 Each reward function captures domain-relevant correctness criteria, assessing whether $o$ satisfies symbolic constraints, passes execution tests, or is preferred over alternatives under subjective evaluation.
@@
 Each task employs a tailored correctness criterion:
-  • Math:
-    $r_{\texttt{math}}(o,a)=\mathbbm{1}\left\{\texttt{verify}_{\texttt{math}}(o_{\texttt{ans}},a)=\texttt{true}\right\}\,,$
-    where $a$ denotes the ground-truth answer, and $o_{\texttt{ans}}$ is the predicted answer extracted from the <answer> tags in the model output $o$.
-Symbolic equivalence between $o_{\texttt{ans}}$ and $a$ is verified using a deterministic parser.
-  • Code Generation:
-[…]
-    where the generated code $o_{\texttt{ans}}$ is executed in a sandboxed environment and evaluated against the unit tests defined by $x$; $\models$ indicates logical satisfaction.
-  • General QA:
-    $r_{\texttt{qa}}(o,a)=\mathbbm{1}\left\{o_{\texttt{ans}}=a\right\}$
+  • Math: Elements of the math dataset are couples $(q,a_{q})$ where $q$ is a prompt and $a_{q}$ is a token sequence stating the answer.
+We implement a $\texttt{verify}_{\texttt{math}}(o,a)$ function that combines regular expression and symbolic parser to checks that $a$ (or an equivalent acceptable answer) is in $o$ within tags <answer>.
+More formally
+    $r_{\texttt{math}}(q,o)=\mathbbm{1}\left\{\texttt{verify}_{\texttt{math}}(o,a_{q})=\texttt{true}\right\}\,.$
+  • Code Generation: Each element of the code dataset is a tuple $(q,\texttt{unittest}_{q},\texttt{test\_case}_{q})$ where $q$ is a prompt, $\texttt{unittest}_{q}$ is a unit test function and $\texttt{test\_case}_{q}$ is a set of test cases.
+Given an output $o$ for prompt $q$, the generated code $o_{\texttt{ans}}$ is extracted from the output $o$ using regular expressions, the unit test $\texttt{unittest}_{q}(o_{\texttt{ans}},x)$ is executed in a sandboxed environment for every test case $x\in\texttt{test\_case}_{q}$.
+More formally
+    $r_{\texttt{code}}(q,o)=\prod_{x\in\texttt{test\_case}_{q}}\mathbbm{1}\left\{\texttt{unittest}_{q}(o_{\texttt{ans}},x)\right\}$
+  • General QA: Each element of the dataset for General QA is a couple $(q,a_{q})$ of prompt and answer.
+The reward is defined by extracting the answer $o_{\texttt{ans}}$ from the output $o$ using regular expressions and testing it against the ground truth $a$.
+More formally:
+    $r_{\texttt{qa}}(q,o)=\mathbbm{1}\left\{o_{\texttt{ans}}=a_{q}\right\}$
     which returns 1 if the predicted answer matches the ground-truth string exactly.
-  • Creative Writing:
-    $r_{\text{writing}}(o,o_{ref})=\begin{cases}1.0&\text{if }o\succ o_{ref}\\ 0.5&\text{if }o\sim o_{ref}\\ 0.0&\text{if }o\prec o_{ref}\end{cases}$
-    where $o\succ o_{ref}$ indicates that the model output $o$ is preferred over the reference $o_{ref}$, $o\prec o_{ref}$ means the reference is preferred, and $o\sim o_{ref}$ denotes a tie.
-Preferences are judged via pairwise comparison by a fixed LLM-as-a-Judge model.
+  • Creative Writing: Each element of the dataset for Creative writing is a couple $(q,o_{\texttt{ref},q})$.
+Given an output $o$ on prompt $q$, the reward is computed by calling a fixed LLM-as-a-Judge model prompted to do a pairwise comparison between $o$ and $o_{\texttt{ref},q}$.
+More formally:
+    $r_{\text{writing}}(q,o)=\begin{cases}1.0&\text{if }o\succ_{q}o_{\texttt{ref},q}\\ 0.5&\text{if }o\sim_{q}o_{\texttt{ref},q}\\ 0.0&\text{if }o\prec_{q}o_{\texttt{ref},q}\end{cases}$
+    where $A\succ_{q}B$ means that the fixed LLM-as-a-Judge model prefered $A$ over $B$ for request $q$, and $A\sim_{q}B$ means it judges the answers as tied.
 Auxiliary Rewards.
 To encourage structured outputs, we define formatting-based rewards shared across tasks:
-$r_{\texttt{format}}(o)=\mathbbm{1}\left\{\texttt{tags\_valid}(o)\right\}$
-$r_{\texttt{tags}}(o)=\frac{1}{4}\cdot\left|\texttt{tags\_present}(o)\right|$
+$r_{\texttt{format}}(q,o)=\mathbbm{1}\left\{\texttt{tags\_valid}(o)\right\}$
+$r_{\texttt{tags}}(q,o)=\frac{1}{4}\cdot\left|\texttt{tags\_present}(o)\right|$
 Here, tags_valid ensures proper nesting of <think> and <answer> tags, while tags_present counts required structural markers.
@@
 Given a model output $o$ and its associated evaluation context $\phi_{k}$, the total reward is computed as:
-$\mathcal{R}_{k}(o,\phi_{k})=\sum_{r\in\mathcal{F}_{k}}w_{r}\cdot r(o,\phi_{k}),$
-where $w_{r}\in[0,1]$ denotes the task-specific weight for reward component $r$.
+$R_{k}(q,o)=\sum_{r\in\mathcal{F}_{k}}w_{r}\cdot r(q,o),$
+where $w_{r}\in[0,1]$ denotes the task-specific weight for the reward component $r$.
 If a component reward is undefined, e.g., due to malformed or unparsable output, it is omitted from the sum.
@@
 A.2 Results
-Figure 5: Backward transfer (BWT) matrix.
-Each cell shows the BWT for a target task (columns) when training a single-task GRPO model on a source task (rows).
-Positive BWT indicates forgetting; negative indicates positive transfer.
-Table 1: Ablation study comparing curriculum Learning (Curr) to its reversed task ordering (Reverse-Curr).
-Ordering tasks from structured to open-ended domains consistently improves performance across all, while reversing this order leads to degraded results, especially in open-ended domains.
-Task Base Model Curr Reverse-Curr
-Math 55.9 59.6 58.2
-General QA 41.5 52.2 22.6
-Code Generation 29.6 35.4 32.7
-Creative Writing 74.2 84.2 75.0
-Table 2: Performance across benchmarks.
-MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math).
-Model Merge applies TIES-merging across tasks.
-Bolded values mark the best per row.
-Domains include Math (7 sets), MMLU-Pro (9 categories), Coding (2 sets), and Creative Writing (MT-Bench).
-Task Base Model ST: Coding ST: Math ST: QA ST: Writing SFT Model Merging Joint MT Curriculum MT
-Math
-aime24 15.33 15.33 14.67 14.00 13.33 16.67 14.67 19.33 14.67
-amc23 62.50 60.00 60.00 62.50 59.00 62.00 65.50 64.00 65.50
-gaokao2023en 72.99 74.29 76.10 74.03 75.58 74.29 74.81 76.62 77.14
-math500 78.20 78.80 80.40 75.40 79.20 76.80 79.80 77.60 81.00
-minerva_math 64.34 63.97 66.54 63.24 61.76 65.07 66.18 68.38 71.69
-OlympiadBench 42.07 42.96 43.70 41.33 42.96 42.96 41.78 43.56 47.41
-General QA
-Biology 57.60 56.76 52.30 67.36 59.00 66.25 65.55 67.22 68.76
-Business 33.46 39.04 25.60 58.68 32.95 48.16 59.82 49.81 47.53
-Chemistry 35.78 31.80 27.30 47.70 38.34 44.08 42.49 42.05 50.71
-Computer Science 53.66 48.05 50.24 55.12 51.95 53.66 53.90 58.78 59.27
-Economics 42.65 49.17 38.74 62.91 44.91 59.60 61.97 56.75 62.09
-Engineering 28.28 31.27 20.43 37.46 26.63 37.77 38.08 35.81 37.05
-Health 46.70 46.21 45.23 50.98 47.19 45.72 52.69 50.73 57.09
-History 37.27 33.33 34.65 47.24 38.58 33.86 47.24 43.31 45.67
-Law 23.16 23.98 20.62 27.88 23.25 26.79 26.61 27.52 29.70
-Math 55.37 52.63 50.41 59.29 56.25 57.36 58.25 59.22 61.21
-Other 44.26 40.04 39.72 50.97 43.94 46.43 51.84 49.89 53.25
-Philosophy 36.87 34.27 33.27 43.89 35.47 38.20 41.48 42.08 42.89
-Physics 41.11 37.41 30.79 53.66 41.57 49.81 46.73 48.04 55.58
-Psychology 50.88 51.50 45.36 60.15 51.75 59.02 59.40 59.27 61.78
-Code Generation
-BigCodeBench 46.49 50.35 46.66 47.10 46.84 44.47 48.07 47.19 49.47
-LiveCodeBench 12.71 21.80 13.07 13.79 13.31 14.21 20.78 17.04 21.30
-Creative Writing
-MT-Bench (Writing) 74.16 71.60 74.16 63.00 78.33 44.17 67.50 83.00 84.17
+A.2.1 Output Format Matters: Full-Text Answers Enhance Generalization
+We examine how output format impacts generalization by comparing models trained to generate full-text answers versus selecting letter choices in multiple-choice QA (MCQ).
+Using GRPO, we train two single-task policies on the training set, one prompted to produce full-text final answers at the end of its chain-of-thought completions, and the other to output only letter choices (e.g., “A”, “B”, “C”).
+Figure 3: Models trained to generate full-text answers outperform those trained to select letter choices. This format promotes deeper semantic understanding over shallow pattern matching or guessing.
+As shown in Figure 3, the model trained to output full-text answers achieves significantly better generalization when evaluated with free-form QA prompts on MMLU Pro (51% vs.
+41%).
+While the letter-choice model slightly outperforms when evaluated strictly on MCQ prompts, the full-text model remains competitive across both prompt formats.
+These results suggest that training with complete, semantically grounded answers encourages deeper reasoning, improving the model’s ability to generalize beyond the specific format seen during training.
+In contrast, letter-choice training risks overfitting to shallow pattern matching, reducing transferability to realistic QA settings that often require articulated responses.
+A.2.2 Curriculum Accuracies Predictions
+On Figure 4 are provided BWT matrices for validation sets and test sets.
+We compute the ranking of curricula using both, see table 4.
+We observe that the Forgettability heuristic provides the second best prediction on test BWT, but exact LOM yields a potentially better candidate.
+Using validation BWT, the curriculum Code-Math-QA-Writing comes second,, but the proposed exact LOM should be rejected based on our entropy discussion.
+Figure 4: Comparison of backward transfer on validation set (left) and test set (right).
+Table 4: Curricula ranked by final geometrical average relative improvement $\overline{\Delta_{\mathrm{rel}}}:=\frac{1}{K}\log{\prod_{k=1}^{K}\frac{\mathrm{Acc}(\theta_{s_{\mathrm{max}}},T_{k})}{\mathrm{Acc}(\theta_{0},T_{k})}}$ on test set (left block) and eval set (right block), with predicted accuracies (%) for each task at each order, and their ranking scores.
+Test set order Validation set order
+Order M C Q W $\overline{\Delta_{\mathrm{rel}}}$ Order M C Q W $\overline{\Delta_{\mathrm{rel}}}$
+$\mathcal{M}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{W}$ 55.2 37.7 51.9 78.4 12.85 $\mathcal{C}\,\mathcal{W}\,\mathcal{M}\,\mathcal{Q}$ 34.9 42.5 26.8 35.4 26.00
+$\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ 57.3 35.4 51.9 78.4 12.22 $\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ 36.7 42.5 27.5 32.6 25.85
+$\mathcal{M}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{W}$ 55.2 36.7 51.0 78.4 11.70 $\mathcal{C}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{M}$ 33.1 42.5 28.1 35.3 25.81
+$\mathcal{M}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{C}$ 55.2 36.1 51.0 75.7 10.43 $\mathcal{W}\,\mathcal{C}\,\mathcal{M}\,\mathcal{Q}$ 34.9 44.7 26.8 33.4 25.77
+$\mathcal{M}\,\mathcal{C}\,\mathcal{W}\,\mathcal{Q}$ 55.2 37.7 51.3 66.6 8.49 $\mathcal{C}\,\mathcal{M}\,\mathcal{W}\,\mathcal{Q}$ 36.7 42.5 26.8 32.8 25.29
+$\mathcal{C}\,\mathcal{M}\,\mathcal{W}\,\mathcal{Q}$ 57.3 35.4 51.3 66.6 7.86 $\mathcal{C}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{M}$ 33.1 42.5 27.4 35.4 25.25
+$\mathcal{M}\,\mathcal{W}\,\mathcal{C}\,\mathcal{Q}$ 55.2 37.1 51.3 64.3 7.22 $\mathcal{C}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{W}$ 34.8 42.5 28.1 32.6 25.09
+$\mathcal{C}\,\mathcal{W}\,\mathcal{M}\,\mathcal{Q}$ 57.2 35.4 51.3 64.3 6.96 $\mathcal{W}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{M}$ 33.1 44.7 27.4 33.4 25.02
+$\mathcal{W}\,\mathcal{M}\,\mathcal{C}\,\mathcal{Q}$ 55.1 37.1 51.3 62.1 6.32 $\mathcal{M}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{W}$ 37.0 40.2 27.5 32.6 24.71
+$\mathcal{M}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{C}$ 55.2 36.1 50.4 64.3 6.07 $\mathcal{W}\,\mathcal{M}\,\mathcal{C}\,\mathcal{Q}$ 35.3 42.3 26.8 33.4 24.64
+$\mathcal{W}\,\mathcal{C}\,\mathcal{M}\,\mathcal{Q}$ 57.2 34.8 51.3 62.1 5.69 $\mathcal{M}\,\mathcal{C}\,\mathcal{W}\,\mathcal{Q}$ 37.0 40.2 26.8 32.8 24.16
+$\mathcal{C}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{W}$ 57.7 35.4 39.5 78.4 5.56 $\mathcal{M}\,\mathcal{W}\,\mathcal{C}\,\mathcal{Q}$ 37.0 42.3 26.8 30.9 23.93
+$\mathcal{W}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{C}$ 55.1 36.1 50.4 62.1 5.18 $\mathcal{Q}\,\mathcal{C}\,\mathcal{W}\,\mathcal{M}$ 33.1 41.3 26.1 35.3 23.23
+$\mathcal{Q}\,\mathcal{M}\,\mathcal{C}\,\mathcal{W}$ 55.6 36.7 38.8 78.4 5.05 $\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ 33.1 43.4 26.1 33.3 23.00
+$\mathcal{C}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{M}$ 57.6 35.4 39.5 75.7 4.67 $\mathcal{Q}\,\mathcal{C}\,\mathcal{M}\,\mathcal{W}$ 34.8 41.3 26.1 32.6 22.52
+$\mathcal{Q}\,\mathcal{C}\,\mathcal{M}\,\mathcal{W}$ 57.7 34.4 38.8 78.4 4.41 $\mathcal{W}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{M}$ 33.1 43.4 25.4 33.4 22.45
+$\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ 55.6 36.1 38.8 75.7 3.77 $\mathcal{M}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{W}$ 37.0 39.1 25.5 32.6 22.14
+$\mathcal{Q}\,\mathcal{C}\,\mathcal{W}\,\mathcal{M}$ 57.6 34.4 38.8 75.7 3.52 $\mathcal{W}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{C}$ 35.3 41.1 24.9 33.4 22.07
+$\mathcal{Q}\,\mathcal{W}\,\mathcal{M}\,\mathcal{C}$ 55.5 36.1 38.8 73.1 2.88 $\mathcal{M}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{C}$ 37.0 41.1 25.5 30.8 21.91
+$\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ 57.6 33.9 38.8 73.1 2.25 $\mathcal{Q}\,\mathcal{W}\,\mathcal{M}\,\mathcal{C}$ 33.5 41.1 26.1 33.3 21.87
+$\mathcal{C}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{M}$ 57.6 35.4 39.0 64.3 0.31 $\mathcal{Q}\,\mathcal{M}\,\mathcal{C}\,\mathcal{W}$ 35.2 39.1 26.1 32.6 21.39
+$\mathcal{W}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{M}$ 57.6 34.8 39.0 62.1 -0.97 $\mathcal{M}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{C}$ 37.0 41.1 24.9 30.9 21.35
+$\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ 55.5 36.1 38.4 62.1 -1.48 $\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ 33.5 41.1 25.4 33.4 21.31
+$\mathcal{W}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{M}$ 57.6 33.9 38.4 62.1 -2.11 $\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ 35.2 41.1 26.1 30.8 21.16
 A.3 Detailed Hyper-Parameters
-We show the hyperparameters used in our training in Table3
-Table 3: Training Hyperparameters for All Training Settings. MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math).
-Hyperparameter Curr. MT Joint MT ST:Coding ST:Math ST:QA ST:Writing SFT
+We summarize the hyperparameters used in our experiments in Table 5. These values were chosen through a combination of prior work, small-scale ablations, and practical compute considerations.
+Table 5: Training Hyperparameters for All Training Settings. ST = Single-Task RL (e.g., ST Math = RL trained only on math).
+Hyperparameter Curr. Learning Joint Training ST Coding ST Math ST QA ST Writing SFT
 Model Configuration
@@
 Training Settings
-   Train Batch Size 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 128
+   Train Batch Size $\times6 25$$\times6$ $\times6 25$$\times6$ $\times6 25$$\times6$ 128
    Learning Rate 1e-6 1e-6 1e-6 1e-6 1e-6 1e-6 2.5e-6
@@
 RL Settings
-   KL Beta 0.00 0.02 0.001 0.04 0.04 0.00 -
+   KL Beta 0.0 0.0 0.0 0.0 0.0 0.0 -
    Clip Ratio Low 0.2 0.2 0.2 0.2 0.2 0.2 -
    Clip Ratio High 0.2 0.2 0.2 0.2 0.2 0.2 -
-   $N$ Rollouts 16 16 16 16 16 16 -
+   NN Rollouts 16 16 16 16 16 16 -
    Rollout Temperature 1.0 1.0 1.0 1.0 1.0 1.0 -